### v1

# On the Fragility of Multimodal Perception to Temporal Misalignment in Autonomous Driving

Abstract
Multimodal fusion (MMF) plays a critical role in the perception of autonomous driving, which primarily fuses camera and LiDAR streams for a comprehensive and efficient scene understanding. However, its strict reliance on precise temporal synchronization exposes it to new vulnerabilities. In this paper, we introduce DejaVu, a novel attack that exploits network-induced delays to create subtle temporal misalignments across sensor streams, severely degrading downstream MMF-based perception tasks. Our comprehensive attack analysis across different models and datasets reveals these sensors’ task-specific imbalanced sensitivities: object detection is overly dependent on LiDAR inputs, while object tracking is highly reliant on the camera inputs. Consequently, with a single-frame LiDAR delay, an attacker can reduce the car detection mAP by up to 88.5%, while with a three-frame camera delay, multiple object tracking accuracy (MOTA) for car drops by 73%. To detect such attacks, we propose Aion, a defense patch that can work alongside the existing perception model to monitor temporal alignment through cross-modal temporal consistency. Aion leverages multimodal shared representation learning and dynamic time warping to determine the path of temporal alignment and calculate anomaly scores based on the alignment. Our thorough evaluation of Aion shows it achieves AUROC scores of 0.92–0.98 with low false positives across datasets and model architectures, demonstrating it as a robust and generalized defense against the temporal misalignment attacks.
I Introduction
Autonomous driving (AD) is designed to navigate and interact with complex environments—a capability fundamentally reliant on a comprehensive understanding of its surroundings. The integration of heterogeneous sensors, such as cameras, LiDARs, and radar, that capture data from different modalities allows an accurate perception with enhanced accuracy [1]. Each individual modality has unique strengths; for example, cameras capture rich semantic details, LiDAR provides accurate depth measurements, and radar particularly excels in detecting speed, even in adverse weather conditions [2]. However, these sensors also face inherent limitations, such as a camera’s sensitivity to lighting variations, LiDAR’s lack of texture information, and radar’s sparsity, which can compromise performance when used independently [3, 4, 5]. On the other hand, multimodal fusion (MMF) refers to the process of integrating such heterogeneous unimodal sensor data to generate a single and comprehensive representation of the surroundings for a more accurate perception [6], ensuring efficient downstream tasks such as object detection, tracking, and segmentation. MMF not only compensates for individual sensor weaknesses but also enhances the robustness of the overall system—a critical assumption that underpins its effectiveness in AD applications [7].
Although MMF has demonstrated considerable effectiveness, it remains a complex problem and a fundamental research challenge due to the intrinsic heterogeneity of sensing modalities, due to unique data formats, resolutions, and noise characteristics [8, 9]. This challenge is further compounded by the critical requirement of temporal alignment during sensor fusion, which is complicated by varying sampling rates across sensors, stemming from underlying hardware limitations [10]. In safety-critical cyber-physical systems like AD, it is obligatory for real-time perception, decision-making, and control. On the other hand, temporal misalignments in sensor data acquisition—such as unexpected delays—can lead to significant degradation in the perception tasks, manifesting as object misdetection, inaccurate localization, and hence, scene misinterpretation. These perceptual inaccuracies undermine the reliability and safety of AD, adversely affecting core functionalities such as longitudinal and lateral control, maneuver execution, safety-critical interventions, etc. [11].
These AD functions demand a perfectly synchronized sensor network with low latency to make timely control decisions [12]. In that regard, Automotive Ethernet (AE) has emerged as a key networking technology for in-vehicular networks, offering data rates from 10 Mbps to 10 Gbps [13]. Moreover, equipped with Time-Sensitive Networking (TSN), which uses generalized Precision Time Protocol (gPTP) for precise clock synchronization, ensures deterministic and real-time communication with bounded latency, which is essential for sensor fusion and control in AD applications [14, 15].
Figure 1: Impact of DejaVu attack on 3D object detection. (a-b) show two benign scenarios without any temporal misalignment, and hence, accurate object detection. (c-d) Illustrate different temporal alignments between camera and LiDAR inputs and highlight how delayed sensor data can lead to incorrect detections, either by detecting non-existent objects (false positives) or missing present ones (false negatives). In (c), the MMF prioritizes the (delayed) LiDAR data and predicts three objects, including a pedestrian who is not present in the current camera view, resulting in a false detection from the camera’s perspective. However, in (d), the MMF still prioritizes the (updated) LiDAR data and predicts two objects, excluding the pedestrian who is present in the current camera view, resulting in a missed detection from the camera’s perspective. In both temporal misalignment attack cases, MMF biased its fusion toward LiDAR, failing to account for semantic discrepancies in the camera modality.
Although the adoption of AE ensures low latency and clock synchronization, temporal misalignment can still occur when data is captured at different sensors at different time points (timestamps), especially due to their varying sampling rates. Some industry-grade software stacks (such as Autoware [^1] https://github.com/autowarefoundation/autoware and Apollo [^2] https://github.com/ApolloAuto/apollo) process the unimodal data independently and use post-processing techniques, such as Kalman filters, to ensure temporal alignment during the fusion based on their corresponding timestamps. Therefore, the efficacy of MMF-based perception relies on two fundamental assumptions. First, $A_{1}$: AE with TSN is secure, which ensures the sensor networks are clock-synchronized and have low latency. Second, $A_{2}$: the timestamp attached to the packet is trustworthy and can be used during the fusion process. However, in this work we propose that an attacker can exploit at least any of $A_{1}$ or $A_{2}$ to compromise the sensor network and expose the MMF to temporal misalignment.
MMF often relies on the assumption that integrating diverse sensor inputs will compensate for individual weaknesses and make it robust against any unimodal attack due to the redundancy of the collected data [8]. However, research shows this redundancy does not inherently provide superior robustness. Both unimodal and multimodal systems can be vulnerable to carefully designed unimodal adversarial or patch attacks [16, 17, 18, 19, 20, 21]. Moreover, research has shown that existing MMF-based perception models lack temporal robustness—the ability to produce reliable output despite temporal inconsistencies—against temporal misalignment and are quite fragile against a tiny delay, even in just one modality [10, 22, 23].
Therefore, to overcome the high overhead of the adversarial or patch attacks but still jeopardize the perception of MMF, we propose a network-based temporal misalignment attack named DejaVu. Under the DejaVu attack, an attacker compromises the in-vehicle network and maliciously creates a misaligned temporal data stream to cause an illusion during perception, even without crafting data. The attacker can achieve this by breaking any of the assumptions that the reliability of MMF is based upon. First, the attacker can break $A_{1}$ by exploiting the existing vulnerabilities in gPTP, and desynchronize sensor networks. PTP has been shown to be vulnerable to various cyberattacks, including man-in-the-middle (MITM) attacks [24], packet delay attacks [25], and spoofing [26, 27]. An attacker who gains such access to the in-vehicular network can manipulate PTP messages by introducing artificial delays, altering timestamps, or replaying old synchronization packets. This can cause different sensors to perceive capture data at slightly misaligned times, leading to misaligned sensor fusion.
Similarly, the attacker can break $A_{2}$ by exploiting the known vulnerabilities in network infrastructure to introduce intensional delays or distortions in the timestamps, causing temporal misalignment in the pre-processing phase of MMF. Particularly, automotive networks typically use UDP [13], a protocol inherently susceptible to man-in-the-middle (MITM) attacks, where attackers can intercept and manipulate communication flows and timestamps without accessing the encrypted data to exacerbate the desynchronization [28, 29]. In both cases, the attacker will be able to break the MMF assumptions of synchronized sensors or trustworthy timestamps, plummeting the perception accuracy. Fig. 1 illustrates how temporal delays due to DejaVu attacks in one of the modalities can degrade the performance of the MVXNet model [30]—a MMF-based 3D object detection model, potentially leading to unsafe driving conditions in AD.
The existing defense techniques against temporal misalignment assume benign timestamps and create alignment based on that. However, they will lack effectiveness against DejaVu, as the timestamps can be compromised as well. Hence, we propose a defense against DejaVu attack by extending this paradigm to include data content analysis along with the timestamps. In that regard, we propose Aion, a defense patch based on multimodal shared representation learning and dynamic time warping to detect and analyze the cross-modal temporal consistency. Aion works as a patch to any existing MMF models, and capitalizes on the semantic coherence of sensor data pairs, enabling the detection of inconsistencies that may arise from malicious temporal desynchronization. In summary, we make the following key contributions:
  • We propose DejaVu, a novel temporal misalignment attack against MMF in AD that exploits vulnerabilities of in-vehicle networks and the fragility of multimodal fusion by selectively delaying sensor streams to disrupt perception in safety-critical tasks.
  • We conduct a comprehensive empirical evaluation of DejaVu on state-of-the-art 3D object detection models (MVXNet [30], BEVFusion [31]) and a multi-object tracking model (MMF-JDT [32]) under various misalignment scenarios using the KITTI [33] and nuScenes [34] datasets. Our findings reveal distinct modality-specific vulnerabilities: object detectors are highly sensitive to LiDAR delays, while the tracking model is significantly impacted by camera timing disruptions. A single-frame LiDAR delay reduces 3D detection mAP by up to 88.5%, and a three-frame camera delay drops multiple object tracking accuracy (MOTA) by 73%.
  • To mitigate these attacks, we introduce Aion, a defense patch that leverages multimodal shared representation learning and dynamic time warping to detect temporal inconsistencies. Aion achieves AUROC scores of 0.92–0.98 with a low false positive rate, demonstrating robust and effective detection of temporal misalignment.
II Related Work
II-A Attacks against MMF
The attacks against MMF can be classified into the following two types:
Sensor Spoofing Attacks.  Sensor spoofing attacks focus on feeding false data to a vehicle’s sensors to mislead the autonomous system. For instance, LiDAR sensors can be fooled by malicious laser signals that fabricate or erase objects in the sensed environment [18]. Beyond lasers, researchers have even created electromagnetic interference to tamper with LiDAR perception, achieving fake point insertions or deletions  [35]. Camera sensors are likewise vulnerable; strategically placed adversarial patches on road signs can cause a ‘Stop sign’ to be ignored or misclassified as another sign [36]. Such single-sensor spoofing attacks can trigger hazardous behaviors – e.g., a phantom obstacle can induce emergency braking, or a hidden obstacle can go unnoticed – underscoring how an attacker controlling one input modality can compromise the whole system.
Such unimodal sensor spoofing attacks can be launched against multimodal model, while the remaining modalities remain unattacked. For example, the attacker can inject fake LiDAR points to create fake objects, aligning them to the camera’s view frustum [19], so it remains consistent with the camera image to avoid inconsistencies during the sensor fusion. On the other hand, it was assumed that an adversarial attack against both sensors at once is unlikely to be successful against AD perception that has multimodal inputs such as camera and LiDAR data. However, researchers have devised attacks that break this assumption by crafting inputs that confound multiple modalities simultaneously. It is possible to generate an adversarial 3D physical object that appears innocuous to humans but fools the fused camera+LiDAR detection pipeline [37]. Attackers have even targeted high-level multi-modal representations by perturbing an image or audio input to make its fused embedding mimic an arbitrary target [38]. These sensor spoofing attacks demonstrate that simply having multiple sensors is not sufficient. With carefully constructed inputs, an adversary can simultaneously mislead camera and LiDAR sensors, defeating the very redundancy meant to ensure safety.
Temporal Delay Attack.  While time delay attacks have not been extensively studied within AD, they have created significant attention in other cyber-physical systems (CPS) domains, such as power systems, wireless networks, unmanned aerial vehicles (UAVs), robotics, and time-sensitive networks (TSNs). In power systems, the PTP is crucial for synchronization measurements and events. However, its vulnerability to time delay attacks poses risks to grid stability [39, 40]. Wireless networks, particularly those used in industrial control systems, are susceptible to time delay attacks that can disrupt communication and control processes [41]. In the realm of UAVs, time delay attacks can significantly impact flight control and mission execution [42, 43]. Moreover, multimodal temporal misalignment in sensor fusion has been shown to degrade the accuracy of simultaneous localization and mapping (SLAM) [44]. Software timing interference can also be exploited to cause system destabilization in CPS [45]. Time-delay attacks also impact the traffic transmission quality under different combinations of Time-Aware Shaper (TAS) or Cyclic Queue Forwarding (CQF) in TSN [46]. Furthermore, vulnerabilities in AE can allow for command injection attacks, leading to replay feed in camera [47]. However, none of the existing works consider the impact of temporal misalignment against task-agnostic MMF-based perception.
Figure 2: Overview of the proposed system modeling with DejaVu attack.
II-B Defense Strategies
In the realm of autonomous vehicle security, various defense mechanisms have been proposed to counteract sensor spoofing and multimodal fusion attacks. These defenses can be broadly categorized into spatiotemporal consistency checks, specification-aware recovery strategies, and hardware-based techniques [48]. One prominent approach involves detecting inconsistencies across different sensor modalities over time. In this defense mechanism, spatiotemporal consistency is checked to identify anomalies in sensor data, such as discrepancies in object motion or position between LiDAR and camera inputs [49]. On the hardware side, manufacturers are improving sensor resilience: newer LiDAR models employ techniques such as pulse signal authentication or unique scanning patterns to make spoofed echoes easier to spot [20]. However, none of these provide the means to detect temporal misalignment attacks in multimodal fusion for AD.
III Temporal Misalignment Attack : DejaVu
This section elaborates on the system model, provides details of the proposed temporal misalignment attacks, and the threat model. Table II in the Appendix contains the summary of notations used throughout the paper.
III-A System Modeling
As shown in Fig. 2, we consider an MMF-based perception system with $m$ sensors denoted as $S_{1},S_{2},\ldots,S_{m}$. At any time $t_{act}$, the transmitted sensor data from $S_{i}$ is represented by $(x_{i}^{(t_{act})},t_{pre})$.
Definition 1 ($x_{i}^{(t_{\text{act}})},t_{\text{pre}}$).
The tuple $(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$ represents a sensor data packet where $x_{i}^{(t_{\text{act}})}$ denotes the data “actually” captured at time $t_{\text{act}}$ (as indicated by the superscript), and $t_{\text{pre}}$ is the associated timestamp metadata, indicating the time at which the sample is “presumed” to have been captured. Discrepancies between $t_{\text{act}}$ and $t_{\text{pre}}$ suggest temporal misalignment.
In MMF-based techniques timestamp $t_{pre}$ is used to ensure proper synchronization before fusing them and under a benign scenario at time $t_{i}$, the sensor data packet data can be expressed as $(x_{i}^{(t_{i})},t_{i})$, where $t_{i}=t_{act}=t_{pre}$. As each of the heterogeneous sensors generates data samples with a unique format and dimension, to ensure an effective sensor fusion, a modality-specific unimodal feature encoder (UFE) $E_{i}$ converts the raw data $x_{i}^{(t_{i})}$ to an intermediate representation $f_{i}^{(t_{i})}=E_{i}(x_{i}^{(t_{i})})$. Such encoders map the raw data to a similar feature format, which allows different sensor fusion techniques, such as straightforward operations like concatenation, merging, average polling, etc., or even more advanced techniques such as tensor-based fusion [50]. Assume that sensor fusion is performed at a global time $t_{\text{sys}}$, using a fusion function denoted by $\mathcal{F}_{\theta}$. This process yields a fused representation $f_{mm}^{(t_{\text{sys}})}$, computed as follows:
$f_{mm}^{(t_{\text{sys}})}=\mathcal{F}_{\theta}\big{(}f_{1}^{(t_{1})},f_{2}^{(t_{2})},\ldots,f_{m}^{(t_{m})}\big{)}$
Finally, the corresponding perception head for the downstream tasks $\mathcal{H}_{\theta}$ predicts the perception results from the fused representation as:
$\mathcal{Y}^{(t_{sys})})=\mathcal{H}_{\theta}\big{(}f_{mm}^{(t_{sys})}\big{)}$
As $t_{i}$ is the actual time when the sensor $S_{i}$’s data was captured, and $t_{sys}$ is the time when the sensor data are synchronized and fused at the central ECU, under normal conditions, the $t_{i}$ can be expressed as:
$t_{\text{sys}}=t_{i}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}}$ (1)
where $\delta_{i}^{\text{syn}}$ represents the nominal temporal misalignment (clock drift) from the global time due to the clock drift of $S_{i}$. This delay is sensor-specific and under a perfect clock synchronization as mentioned in $A_{1}$; it becomes zero ($\delta_{i}^{\text{syn}}\approx 0$). On the other hand, $\delta_{i}^{\text{com}}$ is the communication latency, which indicates the time needed for the data to reach the sensor fusion module. With fast communication infrastructure (AE with TSN), this can be eliminated ($\delta_{i}^{\text{com}}\approx 0$). Therefore, under an ideal condition (such as $A_{1}$ $A_{2}$), $t_{sys}\approx t_{1}\approx t_{2}\approx\dots\approx t_{m}$ for $m$ sensors, ensuring that data from all sensors corresponds to the same system time $t_{sys}$.
III-B Proposed Temporal Misalignment Attack: DejaVu
In this part, we elaborate on our proposed temporal misalignment attack, named DejaVu attack, where an adversary maliciously introduces a timing delay $\delta_{i}^{\text{mal}}$ to one or more sensors with a goal of creating misaligned sensor fusion and creating a false perception of the surroundings. Under DejaVu attack, the attacker compromises the network to add an additional malicious delay of $\delta_{i}^{\text{mal}}$. Hence, the attacker transmits a compromised data packet $\big{(}\tilde{x}_{i}^{(\tilde{t}_{i})},t_{i}\big{)}$ to the fusion ECU with the outdated semantic content $\tilde{x}_{i}^{(\tilde{t}_{i})}$ which was captured at global time $\tilde{t}_{i}$ instead of $t_{i}$, but with the updated time stamp $t_{i}$. In this case, $t_{i}=\tilde{t}_{i}+\delta_{i}^{\text{mal}}$, and the attacker delayed that transmission by $\delta_{i}^{\text{mal}}$ either by breaking (such as $A_{1}$ $A_{2}$). Therefore, $t_{sys}$ can be expressed as:
$t_{sys}=\underbrace{\tilde{t}_{i}+\delta_{i}^{\text{mal}}}_{\displaystyle{t_{i}}}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}}$ (2)
The malicious delay $\delta_{i}^{\text{mal}}$ can vary over time–depending on the attacker’s intent, and be crafted to cause maximum disruption in sensor fusion. In a benign scenario, there will be no temporal misalignment, and sensor data will be received and processed in a timely manner. However, the attacker can launch different forms of DejaVu attack by controlling $\delta_{i}^{\text{mal}}$ in different ways. For instance, in Table I, we provide two possible DejaVu attacks scenarios against $S_{i}$, which are elaborated below:
TABLE I: Two Types of DejaVu Attack Strategy Against $S_{i}$
Attack Name Attack Type Delay Distribution $\delta_{i}^{\text{mal}}$
Constant Delay Constant Constant, $k$
Random Delay Random $\text{Uniform}(0,k)$
Constant Delay
This attack strategy introduces a fixed delay of $k$ frames in the sensor data stream. The attacker can achieve this by creating desynchronization among the clocks, tampering with sensor timestamps, or delaying data at the communication layer. Impact: The fusion model still receives temporally consistent data but with a lag in the target modality. This can lead to delayed (misaligned) perception (i.e., missing or shifted bounding boxes) in critical applications like object detection, where real-time perception is essential.
Random Delay
In this attack strategy, each frame experiences a different delay, randomly sampled from the range $[0,k]$. This strategy not only disrupts the real-time requirement but also disrupts the temporal sequence of sensor data, which is crucial for object tracking. Impact: The fusion system struggles to maintain the proper ordering of sensor inputs, leading to degraded perception accuracy. This will cause erratic behavior in time-sensitive sequential applications, particularly for object tracking and autonomous navigation.
III-C Attackers Objective
The adversary’s primary objective is to compromise the integrity and reliability of the sensor fusion by introducing deliberate temporal misalignment in one or more sensors. The attacker aims to disrupt the coherent integration of multi-modal sensor data, thereby inducing errors in perception. The resultant temporal inconsistencies can lead to incorrect perception results, particularly in safety-critical applications such as AD. Hence, the attacker’s key objective is to disrupt the perception to deteriorate the downstream tasks and hence, compromise the safety and integrity of the control decision. Based on the attacker’s capabilities, she can compromise either only one modality, which we define as unimodal DejaVu attack (Uni-DejaVu), or she can compromise multiple modalities together, defined as multi-modal DejaVu attack (Mul-DejaVu). Under the Uni-DejaVu attack with compromised, where only sensor $S_{1}$ is compromised, the compromised perception result $\tilde{\mathcal{Y}}^{(t_{sys})}$ at time $t_{sys}$ can be expressed as:
$\tilde{\mathcal{Y}}^{(t_{sys})}=\mathcal{H}_{\theta}\big{(}\mathcal{F}_{\theta}\big{(}\tilde{f}_{1}^{(\tilde{t}_{1})},f_{2}^{({t}_{2})},\dots,f_{m}^{({t}_{m})}\big{)}\big{)}$ (3)
Similarly, under the Mul-DejaVu attack with all the sensors compromised, the compromised perception result $\tilde{\mathcal{P}}(t_{sys})$ at time $t_{sys}$ can be expressed as:
$\tilde{\mathcal{Y}}^{(t_{sys})}=\mathcal{H}_{\theta}\big{(}\mathcal{F}_{\theta}\big{(}\tilde{f}_{1}^{(\tilde{t}_{1})},\tilde{f}_{2}^{(\tilde{t}_{2})},\dots,\tilde{f}_{m}^{(\tilde{t}_{m})}\big{)}\big{)}$ (4)
Under the DejaVu attack, the system will perceive the surroundings in a different way than that of the benign case $\tilde{\mathcal{Y}}^{(t_{sys})}\neq{\mathcal{Y}}^{(t_{sys})}$, which can lead to erroneous vehicle control decisions.
III-D Attacker Capabilities
We consider a sophisticated adversary capable of gaining access to the in-vehicle network, particularly AE, and launching targeted attacks that introduce temporal misalignment in MMF. We assume the attacker gains in-vehicle network access through one or more of the following realistic entry points [13]: Physical Access: The attacker connects to the vehicle’s OBD-II port or directly to Ethernet/CAN interfaces, either through malicious maintenance personnel or via physical compromise [51]. Remote Exploitation: The attacker exploits vulnerabilities in externally exposed interfaces, such as telematics units, infotainment systems, or over-the-air (OTA) update mechanisms [51, 52]. Supply Chain Attacks: The attacker implants malicious code or hardware during manufacturing, allowing persistent access post-deployment [53]. Once access is established, the attacker can monitor, intercept, and inject messages on the in-vehicle Ethernet backbone and associated sub-networks. This enables manipulation of time-sensitive communications, particularly those involved in clock synchronization and timestamp propagation.
We assume the attacker operates under two distinct capabilities. $C_{1}$ Disruption of Clock Synchronization.  The adversary targets the clock synchronization mechanism in AE—specifically, the gPTP. Rather than altering timestamps directly, the attacker compromises the synchronization process itself, thereby inducing actual temporal misalignment between sensor streams. This can be accomplished by impersonating the grandmaster clock or tampering with synchronization messages via selective delay, replay, or man-in-the-middle (MITM) attacks [27, 24, 54]. Although individual timestamps remain unaltered, they no longer correspond to a consistent global time due to induced drift. Consequently, sensor data arrives with nominally valid timestamps that are, in fact, temporally skewed, impairing the performance of the downstream perception task. $C_{2}$ Manipulation of Timestamp Integrity.  In this scenario, the attacker is capable of preserving the actual timing of data transmission but alters the timestamps embedded in transmitted packets, particularly in UDP-based communication [13]. This creates a seemingly temporal misalignment while data samples are, in fact, temporally aligned. As a result, while the fusion algorithm aligns the seemingly misaligned inputs, in reality, it creates actual temporal misalignment. Thus, we consider an attacker equipped with either $C_{1}$, $C_{2}$, or both capabilities to launch the DejaVu attack.
IV Temporal Misalignment Defense: Aion
Figure 3: Overview of the proposed defense Aion against any temporal misalignment attack.
To defend against such temporal misalignment attacks, we propose a countermeasure technique named Aion, that can detect if any of the sensor data streams are misaligned. We design Aion as an independent detection patch that can work on top of any MMF-based application, either in parallel or sequentially, agnostic of the downstream task. Fig 3 shows the overview of Aion. In the following sections, we present an overview of the defense and each technical detail.
IV-A Aion Overview
As shown in Fig. 3, Aion primarily consists of a single shared multimodal representation encoder (MRE) that maps any unimodal feature representation, regardless of its source or modality, to a shared representation space. Furthermore, Aion has two phases of implementation: i) development and ii) deployment phase.
Development Phase. In the development phase, Aion trains an MRE model using contrastive learning with positive and negative pairs based on their related temporal alignment. MRE learns how to represent temporally aligned (positive) feature pairs from different modalities to similar representations and temporally misaligned (negative) pairs to different representations. Once the MRE is trained, Aion utilizes that trained MRE in the deployment phase to detect DejaVu attacks.
Deployment Phase. During the deployment phase, Aion utilizes the trained MRE to create shared representations of historical inputs from each modality and keeps a stack of these representations for a small window. At the same time, Aion also calculates and keeps track of a 2D similarity matrix with pairwise inter-modality similarity scores between different representation pairs. The diagonal elements in the similarity matrix indicate pairs that are temporally aligned and others that are temporally misaligned to different extents as they deviate from the diagonal. On each such distance matrix, Aion runs a dynamic time warping (DTW) algorithm to find the optimal path of temporal alignment and the reward of such alignment, which is the summation of all their similarity scores. Under a benign scenario, the optimal path with the highest reward would be the diagonal one, and the reward would be higher. However, under a temporal misalignment attack, the optimal path would deviate from the diagonal and follow the attacker’s misaligned pattern. In that case, the optimal reward would be lower, which essentially indicates the existence of an adversary. We elaborate on the details of each component of Aion in the following subsections.
IV-B Technical Details of Aion
To learn a unified representation for multimodal inputs, we use a shared MRE that projects modality-specific features $f_{i}$ and $f_{j}$ from different modalities into a common latent space, such that $r_{i}^{(t_{i})}=E_{mm}(f_{i}^{(t_{i})})$ and $r_{j}^{(t_{j})}=E_{mm}(f_{j}^{(t_{j}})$. The objective is to ensure that the shared representations of semantically corresponding—i.e., temporally aligned—inputs are close in the latent space, meaning $r_{i}^{(t_{i})}=r_{j}^{(t_{i})}$ if $t_{i}=t_{j}$, and dissimilar otherwise. As the majority of MMF–based perception models for AD primarily focus on fusing camera and LiDAR data ($n=2$), we center our technical discussion of Aion on these two modalities. This choice aligns with common practice in the field. Thus, for the remainder of this section, we assume a two-modality setting involving sensors $S_{i}$ and $S_{j}$, which we refer to as $S_{C}$ and $S_{L}$, denoting camera and LiDAR, respectively. Nonetheless, the proposed Aion framework is designed to be modality-agnostic and readily generalizable to scenarios involving more than two sensor modalities.
Moreover, under benign operating conditions, the sensors are assumed to be synchronized, with data from all modalities captured simultaneously at regular intervals. For notational convenience, instead of referring to the temporal sequences $t_{i}$ and $t_{j}$ which are continuous variables we consider their discrete time steps definied as $n_{i}$ and $n_{j}$, where they belong to the indices set $\mathcal{I}_{total}=\{1,2,3,\dots\}$. Thus, both $n_{i}$ and $n_{j}$ refer to the time indices for the corresponding sensors, with $r_{C}^{(n_{i})}$ indicating the shared representation of $(x_{C}^{(t_{i})},t_{i})$ , and $r_{L}^{(n_{j})}$ representing the shared representation of $(x_{L}^{(t_{j})},t_{j})$.
IV-B1 Development Phase
The development phase specifically involves the training of the MRE model and running the detection on benign data to set the threshold. To ensure effective learning, we utilize contrastive learning with three types of data pairs for the model training.
Different Representation Pairs.  We categorize any representation pairs into three types based on their extent of temporal (mis-)alignment.
Definition 2 (Positive Pairs).
A pair $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})$ is considered a positive pair, denoted $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})\in\mathcal{T}_{p}$, if the corresponding sensor inputs are from the same temporal event, i.e., $n_{i}=n_{j}$.
Definition 3 (Near-Negative Pairs).
A pair $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})$ is considered a near-negative pair, denoted $(r_{C}^{(n_{i})},r_{L}^{(n_{j})}\in\mathcal{T}_{nn}$, if the corresponding sensor inputs are from different but temporally adjacent events, such that $n_{i}\neq n_{j}$ but $n_{i}\approx n_{j}$. In these pairs, $r_{C}^{(n_{i})}$ and $r_{L}^{(n_{j})}$ share partially overlapping semantic content due to their temporal proximity.
Definition 4 (Far-Negative Pairs).
A pair $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})$ is considered a far-negative pair, denoted $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})\in\mathcal{T}_{fn}$, if the sensor inputs originate from totally distinct temporal events with no temporal or semantic overlap; that is, $n_{i}\neq n_{j}$ and $n_{i}\not\approx n_{j}$.
Objective Function.  The primary goal of the shared encoder $E_{mm}$ is to ensure that the representations of positive pairs are highly similar—i.e., have minimal distance—while representations of negative pairs remain well separated in the latent space. To achieve this, we adopt a contrastive learning objective, based on relaxed contrastive (ReCo) as proposed in [55], to train $E_{mm}$, where each training batch consists of a set of discrete sample indices $\mathcal{I}_{batch}=\{n_{1},n_{2},\dots,n_{b}\}$, where the batch size is $b$ and each $n_{k}$ corresponds to a unique sample in the batch.
Thus, the representation sequences $\mathbf{r}_{C}=\{r_{C}^{(n_{1})},r_{C}^{(n_{2})},\dots,r_{C}^{(n_{b})}\}$ and $\mathbf{r}_{L}=\{r_{L}^{(n_{1})},r_{L}^{(n_{2})},\dots,r_{L}^{(n_{b})}\}$ from two different modalities are calculated on the sampled inputs from the training set. These indices are chosen in a manner that ensures the batch contains positive, near-negative, and far-negative pairs. Based on the $\mathbf{r}_{C}$ and $\mathbf{r}_{L}$, we compute a similarity matrix $\mathbf{S}\in\mathbb{R}^{b\times b}$, where each entry $S_{ij}$ denotes the cosine similarity between the camera representation $r_{C}^{(n_{i})}$ and the LiDAR representation $r_{L}^{(n_{j})}$, defined as:
$S_{ij}=\frac{r_{C}^{(n_{i})}\cdot r_{L}^{(n_{j})}}{\|r_{C}^{(n_{i})}\|\,\|r_{L}^{(n_{j})}\|}$ (5)
For positive pairs—i.e., pairs with maximum semantic alignment corresponding to the same temporal event—we expect the cosine similarity to approach 1. To enforce such alignment, we define the positive loss as:
$\mathcal{L}_{\text{pos}}=\sum_{i=1}^{b}\left(S_{ii}-1\right)^{2}$ (6)
This loss encourages the cosine similarity between the shared representations of temporally aligned inputs—i.e., $r_{C}^{(n_{i})}$ and $r_{L}^{(n_{i})}$—to be as close as possible to 1.
Negative pairs consist of temporally misaligned inputs—either near-negatives (slightly misaligned) or far-negatives (strongly misaligned). Ideally, their representations should exhibit minimal cosine similarity. To enforce this, we define the negative loss as:
$\mathcal{L}_{\text{neg}}=\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{b}\left(\max(0,S_{ij})\right)^{2}\cdot\lambda_{ij}$ (7)
This loss penalizes cases where misaligned representations $r_{C}^{(n_{i})}$ and $r_{L}^{(n_{j})}$ exhibit undue similarity. The penalty is modulated by the weight $\lambda_{ij}$, which reflects the expected degree of dissimilarity based on temporal distance. For example, we hypothesize the weight $\lambda_{ij}$ based on the type of pair:
$\lambda_{ij}=\begin{cases}\lambda\in(0,1),&\text{if }(r_{C},r_{L})\in\mathcal{T}_{nn}\quad\text{(near-negatives)}\\ 1,&\text{if }(r_{C},r_{L})\in\mathcal{T}_{fn}\quad\text{(far-negatives)}\end{cases}$
To generalize this weighting scheme, we define $\lambda_{ij}$ as a smooth function of temporal distance:
$\lambda_{ij}=\tanh\left(\frac{|n_{i}-n_{j}|}{\tau}\right)$ (8)
Here, $\tau$ is a temperature-like scaling factor that controls sensitivity to temporal separation. This formulation offers a continuous and differentiable measure of misalignment, encouraging the model to learn nuanced distinctions across the temporal spectrum. Intuitively, near-negative pairs may retain some semantic overlap, and are thus penalized less for being similar. In contrast, far-negative pairs—expected to be semantically unrelated—are penalized more heavily if their representations remain close. This enables the model to learn robust, temporally-aware embeddings while maintaining sensitivity to semantic drift. The overall objective combines the positive and negative pair losses:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{neg}}$ (9)
This formulation promotes high cosine similarity for aligned (positive) pairs, while pushing apart misaligned (negative) pairs. The extent of separation for negative pairs is controlled by the penalty weight $\lambda_{ij}$, allowing for flexibility based on temporal misalignment.
Algorithm 1 Optimal Warping Path and Reward Computation
1:  Input: Cost matrix $\mathbf{S}\in\mathbb{R}^{w\times w}$
2:  Output: Optimal warping path $\mathcal{P}^{*}$ and total optimal reward $R(w,w)$
3:  Initialize accumulated similarity matrix $R\in\mathbb{R}^{w\times w}$
4:  Set $R(1,1)=S(1,1)$Initialize first element cost}
5:  for $n=2$ to $w$ do
6:     $R(n,1)=S(n,1)+R(n-1,1)$
7:  end for
8:  for $m=2$ to $w$ do
9:     $R(1,m)=S(1,m)+R(1,m-1)$
10:  end for
11:  for $n=2$ to $w$ do
12:     for $m=2$ to $w$ do
13:        $R(n,m)=S(n,m)+\max\left(R(n-1,m-1),R(n-1,m),R(n,m-1)\right)$
14:     end for
15:  end for
16:  Backtracking to find the optimal warping path:
17:  Initialize $\mathcal{P}^{*}=[(w,w)]$Start from the last position}
18:  Set $(n,m)=(w,w)$
19:  while $(n,m)\neq(1,1)$ do
20:     if $n=1$ then
21:        $m\leftarrow m-1$
22:     else if $m=1$ then
23:        $n\leftarrow n-1$
24:     else
25:        Choose $(n,m)$ from $\arg\max\left(R(n-1,m-1),R(n-1,m),R(n,m-1)\right)$
26:     end if
27:     Prepend $(n,m)$ to $\mathcal{P}^{*}$
28:  end while
29:  reward $\phi^{*}=R(w,w)$
30:  Return: $\mathcal{P}^{*}$ and $\phi^{*}$
IV-C Attack Detection
The detection of DejaVu attacks, though Aion consists of two main tasks. First, Aion maintains individual queues of historical representations for each modality, and a similarity matrix indicating pair-wise similarity. Second, Aion analyzes the similarity matrix using dynamic time warping to find the optimal path of temporal alignment. From the observed optimal path and its total reward, Aion calculates the anomaly scores and detects any potential temporal misalignment. We elaborate on each step in the following subsections.
IV-C1 Historical Representation Queue and Similarity Matrix
Aion keeps queues of historical representations for each modality for the latest $w$ sensor data. If we assume the indices of that queue as $\mathcal{I}_{detect}=\{n_{1},n_{2},\dots,n_{w}\}$. For a presented window size $w$, Aion keeps track of the sequential representation $\mathbf{r}_{C}=\{r_{C}^{(n_{1})},r_{C}^{(n_{2})},\dots,r_{C}^{(n_{w})}\}$ and $\mathbf{r}_{L}=\{r_{L}^{(n_{1})},r_{L}^{(n_{2})},\dots,r_{L}^{(n_{w})}\}$. Similar to the training phase, as mentioned in (5), Aion creates the similarity matrix $\mathbf{S}\in\mathbb{R}^{w\times w}$. With the arrival of every new message, Aion updates $\mathbf{r}_{C}$, $\mathbf{r}_{L}$, and $\mathbf{S}$, and runs the DTW-based detection as described in the following part.
IV-C2 Dynamic Time Warping-based Detection
To quantify the extent of temporal misalignment within the $\mathbf{r}_{C}$ and $\mathbf{r}_{L}$, Aion employs DTW to compute both the optimal temporal alignment path and the corresponding alignment reward. We implemented DTW to identify the optimal warping path $\mathcal{P}$ that maximizes the accumulated similarity, which we define as reward, over a similarity matrix. Algorithm 1 outlines this procedure, which takes $\mathbf{S}$ as input and returns the optimal path $\mathcal{P}$ and total reward $\phi$ associated with that path. In an ideal scenario, where all sensors remain temporally aligned, the optimal warping path follows the diagonal: $\mathcal{P}^{*}=\{(1,1),(2,2),\dots,(w,w)\}$, as diagonal elements $S_{ii}$ have the highest similarity scores. Under the optimal alignment path, the optimal accumulated reward, $\phi^{*}=\sum_{i=1}^{w}S_{ii}=w$, since the embedding function $E_{mm}$ is trained to maximize similarity for aligned pairs. Thus, any deviation from that diagonal path $\mathcal{P}^{*}$ or the optimal reward $\phi^{*}$ can be considered anomalous.
For any test case with optimal path $\mathcal{P}$ and total reward $\phi$, the anomaly score $a$ can be considered as $a=\phi^{*}-\phi=w-\phi$. In a benign scenario, all the sensors remain temporally aligned, and the optimal warping path usually follows the diagonal line ($\mathcal{P}_{ben}\approx\mathcal{P}^{*}$), as diagonal elements $S_{ii}$ have the highest similarity scores. Since $E_{mm}$ is trained to maximize similarity for aligned pairs, resulting in a high alignment reward $\phi_{ben}\approx w$ and low anomaly score $a_{ben}\approx 0$. Conversely, in the presence of a malicious actor in the network triggering temporal misalignment, the optimal warping path $\mathcal{P}_{mal}$ deviates from $\mathcal{P}^{*}$, incorporating non-diagonal elements where $n_{i}\neq n_{j}$, lowering the total reward $\phi_{mal}$ and hence, increasing the anomaly score $a_{mal}$.
Justification on Detection.  The fundamental assumption behind this approach is that DTW maximizes cumulative alignment reward by optimally aligning sequences. Given a well-trained $E_{mm}$, the cost function $S_{ij}$ satisfies:
$S_{ij}=1\quad\text{iff}\quad n_{i}=n_{j}$
In a benign case, where data samples are perfectly aligned, $\phi_{ben}$ is maximized, and $a_{ben}$ is minimized, since all elements on the optimal path mostly satisfy $n_{i}=n_{j}$, therefore:
$\phi_{ben}=\sum_{(i,j)\in\mathcal{P}_{ben}}S_{ij}\approx\sum_{i=1}^{W}S_{ii}\quad\text{thus,}\quad a_{ben}\approx 0$
However, in the presence of malicious misalignment, the warping path necessarily includes terms where $n_{i}\neq n_{j}$, leading to $S_{ij}<<1$ for some $(i,j)$. Since DTW optimally maximizes the total reward, the deviation from $\mathcal{P}^{*}$ implies a decrease in $\phi_{mal}$ and an increase in $a_{mal}$ is minimized, such that:
$\phi_{mal}=\sum_{(i,j)\in\mathcal{P}_{mal}}S_{ij}<\sum_{i=1}^{W}S_{ii}\quad\text{thus,}\quad a_{mal}>>0$
This establishes the fundamental assumption that as misalignment increases, so does anomaly score, reinforcing the validity of DTW in the anomaly detection process. Empirical validation in Section IV-C further supports this claim.
V Evaluation on DejaVu
We present the evaluation results of the DejaVu attack applied to MMF in the context of AD. This section details the experimental setup and discusses the key findings from the evaluation.
V-A Experimental Settings
We describe the experimental setup used to assess the DejaVu attack across different MMF models and datasets.
V-A1 Datasets
We experiment with two widely used multimodal autonomous driving datasets for both 3D object detection and multi-object tracking. KITTI Tracking Dataset. The KITTI Tracking Dataset [33] is a widely used benchmark for object detection and multi-object tracking in AD. It provides RGB images, LiDAR point clouds for a variety of urban, suburban, and highway driving scenarios for multiple object categories, including cars, pedestrians, and cyclists. NuScenes Dataset.  The NuScenes Dataset [34] is a large-scale autonomous driving dataset that includes multi-modal sensor data, including six RGB cameras, a LiDAR with multiple categories, such as vehicles, pedestrians, and barriers.
V-A2 Models
This part elaborates on the state-of-the-art MMF models we consider to test the DejaVu attacks. To systematically evaluate the effect of DejaVu on MMF with different downstream tasks, we consider the following models:
3D Object Detection: We evaluate DejaVu on two representative MMF-based 3D object detection models. i) MVXNet [30], trained on KITTI dataset, is an early fusion-based architecture that projects LiDAR point clouds into pseudo-image space and fuses them with camera image features at the voxel level. ii) BEVFusion [31] is a recent architecture, trained on the NuScenes dataset, unifies multi-modal sensor inputs—including LiDAR and camera—in the bird’s eye view (BEV) representation space. It employs modality-specific encoders followed by a fusion module that generates a shared BEV feature map for downstream detection.
Multi-Object Tracking (MOT): For evaluating tracking performance under DejaVu, we consider MMF-JDT [32], trained on the KITTI tracking dataset, is a joint detection and tracking model that incorporates early and mid-level fusion strategies to align image and point cloud features for improved object association over time.
V-A3 Attack Settings
To evaluate the impact of DejaVu attack on MMF models, we introduce controlled delays in one or both sensor modalities (camera and LiDAR), as mentioned in Table I, and analyze the corresponding degradation in model performance. We systematically assess how different degrees and types of temporal misalignment affect the 3D object detection and multi-object tracking. We introduce both constant and random delays in one or both modalities, as defined in Table I. Specifically, we use a delay parameter $k\in\{0,1,2,3,4,5\}$, where $k=0$ represents a perfectly synchronized sensor, and $k=5$ denotes a maximum delay of five frames for the affected modality. We consider both Uni-DejaVu and Mul-DejaVu attacks. In Uni-DejaVu attack, delay is introduced in either the camera or the LiDAR input while keeping the other modality synchronized. In Mul-DejaVu attacks, both camera and LiDAR streams are delayed independently, leading to varying degrees of temporal misalignment. We use the pretrained weights for the target model provided with the official implementations. We focus exclusively on attacking the test dataset by applying the defined temporal delays.
V-A4 Evaluation Metrics
To assess the impact of DejaVu, we analyze the model performance using task-specific evaluation metrics. For 3D object detection, we evaluate the models using mean Average Precision (mAP), which quantifies the accuracy of detected objects. For MOT, we use standard tracking metrics such as higher order tracking accuracy (HOTA), detection accuracy (DetA), association accuracy (AssA), multiple object tracking accuracy (MOTA), and Identity Switches (IDSW), which measure the effectiveness of object association across frames.
V-B Software Implementation
We implement both DejaVu and Aion using Python 3.8 and PyTorch, utilizing open-source frameworks including MMDetection3D [56] and OpenPCDet [57]. Experiments were conducted on a server running Ubuntu 20.04.6 LTS with an Intel Xeon Gold 5520 (16 cores, 2.20GHz), 128GB RAM, and three NVIDIA RTX 6000 Ada GPUs.
V-C Evaluation Results
This section presents the DejaVu attack impact on different MMF models and datasets.
V-C1 Impact of DejaVu Attack on 3D Object Detection
We investigate the effectiveness of the proposed DejaVu attack on multimodal 3D object detection using MVXNet and BEVFusion, evaluated on the KITTI and nuScenes datasets, respectively. We analyze the detection performance of 3D object detection across different object classes under varying levels of unimodal and multimodal sensor delay. Object detection models do not process sequential information; instead, their performance is affected only by frame-wise delays in each modality at any particular time, whether the delays are constant or random. Therefore, for simplicity and consistency, we only analyze the attack under the constant delay setting.
MVXNet on KITTI Dataset
Fig 4 presents the 3D object detection performance of MVXNet under Uni-DejaVu attacks, where either the camera or LiDAR input is delayed independently. Under benign (zero-delay) conditions, MVXNet achieves strong mAP across most object classes: approximately 84.1 for cars, 87.6 for pedestrians, and 73.2 for cyclists. The left plot shows that delaying the camera input alone–under Uni-DejaVu camera attacks—has minimal effect on performance across all classes, with nearly constant mAPs. In contrast, the right plot highlights the model’s high sensitivity to LiDAR delays: a 1-frame delay causes the car mAP to collapse from 84.1 to 9.7 ($\downarrow$88.5%), pedestrian mAP from 87.6 to 34.2 ($\downarrow$60.9%), and cyclist mAP from 73.2 to 32.9 ($\downarrow$55.1%), with further degradation as delay increases. This indicates that LiDAR data is significantly more critical than camera input in MVXNet’s perception pipeline; hence, MVXNet’s high vulnerability against Uni-DejaVu attack against LiDAR. However, Fig. 5 shows the mAP heatmaps across combinations of camera and LiDAR delays under Mul-DejaVu attacks. Although the 1-frame LiDAR delay drops mAP from 55.1% to 88.5% for different objects, further camera delay along with LiDAR has no additional effect, indicating the dominance of LiDAR in MVXNet.
Key Findings: MVXNet heavily depends on LiDAR input for 3D object detection. Camera delay, under both Uni- or Mul-DejaVu attack, has almost no effect, but even minor LiDAR misalignment leads to severe performance degradation. When both modalities are delayed, the impact is still only dominated by the LiDAR stream.
Figure 4: Uni-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
Figure 5: Mul-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
Figure 6: Uni-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
Figure 7: Mul-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
(a) Constant Delay Scenario
(b) Random Delay Scenario
Figure 8: DejaVu attack impacts on multi-object (car) tracking performance of MMF-JDT on KITTI tracking dataset with respect to different metrics.
BEVFusion on nuScenes Dataset
Fig. 6 shows BEVFusion’s 3D detection performance under Uni-DejaVu attacks. With no delay, the model achieves high mAP across most classes—approximately 88 for car and pedestrian, and slightly lower for the rest. With a 1-frame camera delay, mAP drops slightly across all classes (for instance, car mAP drops by 7.4%), but remains stable with further delays, showing BEVFusion is slightly vulnerable against Uni-DejaVu camera attacks. In contrast, introducing a 1-frame LiDAR delay results in a substantial reduction in mAP for specific object classes, with performance dropping by 65.6% for cars (from 88.8 to 30.5), 35.1% for pedestrians (from 88.1 to 57.2), and 38.6% for motorcycles (from 75.9 to 46.6). Although performance remains stable with further LiDAR delays, this finding shows BEVFusion’s higher vulnerability against Uni-DejaVu LiDAR attacks. Similarly, Fig. 7 illustrates BEVFusion’s performance under Mul-DejaVu attacks. Although delaying the camera alone has a minimal impact on performance, combining this with a delay in the LiDAR stream causes a significant drop in mAP. For instance, a 1-frame delay in camera or LiDAR stream reduces car mAP by 7.4% and 65.5%, respectively. However, when both modalities are delayed simultaneously by one frame, the mAP drops dramatically by 89.2% (from 88.8 to 9.6), highlighting a substantial compounding effect. This trend is consistently observed across all object classes.
Key Findings.  BEVFusion is slightly affected by camera delays but is highly affected by LiDAR delays, further underscoring LiDAR’s dominating role in 3D object detection. However, the impact becomes significant if both sensors are delayed simultaneously, even just by one frame.
V-C2 Impact of DejaVu Attack on Multi Object Tracking
We investigate the effectiveness of the proposed DejaVu attack on MOT algorithm using MMF-JDT on the KITTI dataset. We analyze the tracking performance for cars under varying levels of delays under Uni- and Mul-DejaVu attack scenarios.
MMF-JDT on KITTI Dataset
Fig 8 illustrates the impact of DejaVu attacks on MMF-JDT evaluated on the KITTI tracking dataset. For this evaluation, we consider both types of delays: constant (Fig 8a) and random (Fig 8b). Across both attack scenarios, the tracking performance declines as camera delay increases. Along with other tracking metrics, MOTA and IDSW suffer noticeable degradation as camera delays increase under Uni-DejaVu attacks. For instance, IDSW—a metric that captures identity switches and ideally should be low, increases dramatically under increasing camera delays, underscoring the disruption in tracking consistency caused by DejaVu attacks.
Furthermore, the values of different metrics across the heatmaps of Fig 8a suggest that while camera delay under Uni-DejaVu deteriorates the performance, delaying both the camera and LiDAR by exactly the same delay (i.e., constant delay scenario) under Mul-DejaVu attack diminishes the attack impact and mostly retains the performance. On the other hand, heatmaps in Fig. 8b show that Mul-DejaVu attacks with random delays remain effective as different delays in both modalities break the sequence, making object tracking considerably more difficult. For instance, under the Mul-DejaVu constant attack scenario with a five-frame delay, MOTA decreases by only 1.4%, whereas the random attack scenario results in a 60.5% drop.
Key Findings.  In contrast to 3D object detection tasks, MOT appears to rely more heavily on camera inputs. This may be attributed to the fact that MOT does not require precise 3D bounding boxes; instead, the rich texture information in camera images may offer more effective contrastive representations than sparse point clouds. As a result, DejaVu attacks can substantially impair MOT performance, particularly under Uni-DejaVu (camera delay) and Mul-DejaVu with random delay scenarios.
VI Evaluation of Aion
VI-A Experimental Settings
To evaluate the effectiveness of Aion in detecting DejaVu attacks, we conduct a detection analysis under various attack scenarios. We synthetically generate different degrees of temporal misalignment by perturbing the input sequences in the test data as described in Table I. For two different models trained on two different datasets, we evaluate Aion’s ability to distinguish between normal and misaligned sequences under diverse DejaVu attacks.
Model Architecture of Aion
We implemented Aion for both the KITTI and nuScenes datasets to evaluate its adaptability across different sensor setups and driving scenarios.
Aion on KITTI: For the KITTI dataset, we adopt a straightforward approach by testing with two off-the-shelf, pre-trained image and LiDAR feature encoders. The MRE of Aion is implemented using a simple convolutional neural network (CNN) architecture, featuring two distinct input branches and a shared output branch. For the KITTI dataset, each input sample consists of an RGB image and a corresponding LiDAR point cloud. The RGB input image has the size of $[3,375,1242]$, and is processed using a pretrained ResNet-50 [58] feature encoder, yields image feature maps $f_{C}$ of shape $[2048,12,39]$. Simultaneously, the LiDAR input of size $[k,3]$, where $k$ is the number of points, is voxelized using a pretrained PointPillars [59] encoder, producing LiDAR feature maps $f_{L}$ of shape $[384,248,216]$. Later, we use our proposed $E_{mm}$ to project both $f_{C}$ and $f_{L}$ to a shared space with representations $r_{C}$ and $r_{L}$, respectively. First, both $f_{C}$ and $f_{L}$ are passed through their modality-specific convolutional input branch that reduces their channel dimension to 256, with progressive spatial downsampling, followed by a global average pooling layer, and a shared fully connected projection head to produce 256-dimensional representation vectors for both $r_{C}$ and $r_{L}$.
Figure 9: Similarity scores between Camera and LiDAR representation embeddings under benign and DejaVu attacks between time steps 25 to 35.
Aion on nuScenes: For the nuScenes dataset, we build Aion directly on top of BEVFusion [31] architecture to showcase Aion’s adaptability with a more complex MMF architecture. Due to the complexity of the nuScenes dataset, where each input contains six camera images and a 360 lidar point cloud, we introduce a hybrid architecture for $E_{mm}$, utilizing both CNN and transformer architectures, to learn compact and robust representations. Using BEVFusion’s own individual input feature encoders, we analyze the six images and create a unified BEV feature map $f_{C}$ with size $[64,180,180]$. Point clouds from lidar are also processed and mapped to the feature vector $f_{L}$ with the same dimension using the slightly modified point cloud encoder, as detailed in [31]. Later, $f_{C}$ and $f_{L}$ are fed to our proposed hybrid $E_{mm}$, which has two parts. First, they are passed through a shared convolutional pipeline composed of three CNN layers to have an intermediate embedding with output shape $[256,23,23]$. Later, we adopt a transformer architecture to process the intermediate representation, where the feature maps are reshaped into spatial tokens with $256$-dimensional embeddings, and positional encodings are added. A lightweight transformer encoder with a single self-attention layer enables global spatial reasoning. The sequence is reduced using mean pooling, resulting in $r_{C}$ or $r_{L}$, which are $256$-dimensional shared representations. For both the KITTI and nuScenes datasets, we train the $E_{mm}$ with multimodal contrastive learning, as mentioned in Section IV-B1, and train the models for $100$ and $500$ epochs, respectively, on the training dataset. While training $E_{mm}$ we consider $\tau=3$, which sets the boundary of near- and far-negative sample pairs, during the loss calculation, as mentioned in (7) and (8).
Figure 10: Visualization of anomaly scores generated by Aion under different DejaVu attacks on different modalities. The ‘red’ and ‘green’ shaded areas indicate the time periods with or without DejaVu attacks, respectively. The distinctive anomaly scores at different regions show Aion’s high effectiveness.
Attack Synthesis
For both datasets, we synthetically launch DejaVu attacks on the test data sequences at certain intervals to create malicious test sequences with varying degrees of misalignment. Specifically, we launch each attack scenario outlined in Table I at every $25$ time steps, which persist for the next $k=10$ continuous time steps. This systematic attack injection ensures a dataset containing misaligned samples with different extents of temporal shifts, allowing for a comprehensive evaluation of the defense mechanism.
Anomaly Detection Methodology.
To classify whether an input sequence is malicious, we analyze the cross-modal temporal consistency of multimodal pairs within a defined observation window $w=5$. In this evaluation, we label a window as malicious if at least one of its multimodal pairs contains a misaligned sample. We analyze the anomaly scores using the ROC curve and calculate the area under the ROC curve (AUROC) as the key detection metric.
VI-B Detection Results
We evaluate the performance of Aion across both datasets and model architectures. We begin by illustrating the detection process on the nuScenes dataset, including visualizations of similarity and anomaly scores under different attack types. Finally, we present the ROC curves, along with the AUROC scores, for both datasets.
VI-B1 Visualization of Similarity Matrix
Fig 9 illustrates four different similarity matrices with the Camera and LiDAR representations, $r_{C}$ and $r_{L}$, from time steps 10 to 50 under various DejaVu attacks (launched from 25 to 35), including the benign case. The left-most panel shows the similarity matrix between $r_{C}$ and $r_{L}$ under benign conditions—i.e., with no delay in either modality. As illustrated, the highest similarity scores lie along the diagonal path from (10, 10) to (50, 50), indicating perfect temporal alignment between both modalities. However, the two middle panels depict cases where two types of temporal misalignments are introduced under Uni-DejaVu attacks: one with a constant delay applied to the camera stream, and another with a random delay introduced in the LiDAR stream, both between time steps 25 and 35. In these scenarios, the highest similarity scores diverge from the diagonal beyond time step 25 and only return to the diagonal again around time step 35. These deviations clearly signify temporal misalignments, which Aion leverages to detect such DejaVu attacks.
The right-most panel presents a unique scenario where both modalities are delayed by the same amount (constant delay) under Mul-DejaVu attack. In this case, the similarity scores remain high (and the same) across both diagonal and off-diagonal elements from time steps 25 to 35. Such patterns may emerge under both benign and malicious conditions. For instance, under benign conditions, the vehicle may be stationary without any moving objects in the scene, resulting in temporally consistent features over time. In contrast, an attacker could also replicate this same scene with malicious delay to all of the modalities by the same constant offset, creating a similar similarity matrix. Hence, these unique advanced attack becomes a challenging task just by analyzing the cross-modal alignment similarities. Although Aion, when limited to only the modalities used in MMF, cannot reliably detect such an advanced attack case, incorporating additional data sources—such as inertial measurements (IMU), controller area network (CAN) signals, or other external references—can provide complementary evidence and help detect such advanced attacks. However, as we only rely on the multimodal data in this work, we consider this extension as a future work for Aion.
VI-B2 Detection Performance of Aion
We illustrate the detection performance of Aion from two different perspectives.
Visualization of Anomaly Scores
Fig. 10 illustrates the temporal evolution of anomaly scores, provided by Aion, across different time steps under various attack scenarios. Each shaded region indicates whether the system is operating under benign (green) or malicious (red) conditions, based on the temporal alignment. As shown, Aion consistently produces higher anomaly scores during periods where temporal misalignment is introduced, compared to benign intervals where no such misalignment is present. This clear contrast in anomaly demonstrates the effectiveness of Aion in detecting malicious temporal misalignment induced by DejaVu attacks on different modalities.
ROC Curve with AUROC Scores
Here, we provide quantitative evaluations of Aion’s detection performance on both nuScenes and KITTI datasets. Fig. 11 presents the ROC curves evaluating the performance of the Aion in detecting DejaVu attacks on the nuScenes dataset under various sensor delay scenarios. The evaluation encompasses both constant and random delay injections on camera, lidar, and both modalities. Aion achieves consistently high AUROC scores, ranging from $0.94$ to $0.97$, demonstrating strong detection capabilities. In particular, constant delays on the camera and lidar sensors yield the highest AUC of $0.97$. Even under more challenging conditions, such as random delays affecting both sensors, the Aion maintains an AUC of $0.94$. These results illustrate that the Aion effectively detects temporal misalignment under DejaVu attacks while maintaining low false positive rates.
Figure 11: ROC curve of the Aion against DejaVu attacks on nuScenes dataset.
Figure 12: ROC curve of the Aion against DejaVu attacks on KITTI dataset.
Fig. 12 illustrates the ROC curves corresponding to the DejaVu detection mechanism on the KITTI dataset across identical attack scenarios as considered above. Similar to nuScenes, Aion also demonstrates high efficacy against KITTI dataset, achieving AUROC values up to $0.98$ for constant delay attacks on camera and lidar individually. In the presence of random delay attacks against one or both modalities, the detection performance remains robust, with AUROC scores ranging from $0.92$ to $0.97$. These findings substantiate the generalizability of the Aion across datasets and model architectures, further emphasizing its practical effectiveness. Moreover, the consistently high true positive rates with a low false positive rates underscore the Aion’s reliability in realistic AD environments subjected to DejaVu attacks.
VII Conclusion
This work presents DejaVu, a novel temporal misalignment attack that exploits synchronization vulnerabilities in multimodal perception systems for autonomous driving. Through extensive evaluations on state-of-the-art 3D object detection and multi-object tracking models, we uncover modality-specific vulnerabilities: 3D detection models are predominantly reliant on LiDAR and suffer severe degradation—with up to 88.5% drop in mAP—from even a single-frame LiDAR delay, while MOT models exhibit heightened sensitivity to camera stream disruptions, with MOTA dropping by 73% under just three-frame camera delays. To mitigate these vulnerabilities, we propose Aion, a detection framework based on multimodal shared representation learning and dynamic time warping. Aion consistently achieves AUROC scores of 0.92–0.98 with low false positive rates on both KITTI and nuScenes datasets. These findings highlight the critical need for synchronization-aware design in perception architectures and emphasize the importance of robust temporal consistency checks in safety-critical autonomous systems.
Appendix A Overall Evaluation Results on BEVFusion
Fig. 13 shows the overall impact of Mul-DejaVu on average mAP and NDS score on the nuScenes dataset.
Figure 13: BEVFusion (nuScnenes): Overall 3D object detection mAP under Mul-DejaVu attacks for different object classes regarding i) NDS, ii) mAP
Appendix B Notation Summary
TABLE II: Summary of Notations
Symbol Description
$S_{i}$ $i$-th unimodal sensor
$m$ Total number of sensor modalities
$(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$ Data packet from $S_{i}$ with capture and transmission times
$x_{i}^{(t_{\text{act}})}$ Sensor reading from $S_{i}$ captured at time $t_{\text{act}}$
$t_{\text{pre}}$ Time when $x_{i}^{(t_{\text{act}})}$ is made available or transmitted
$E_{i}$ Encoder for modality $i$ (unimodal feature extractor)
$f_{i}^{(t_{i})}$ Feature representation of input $x_{i}^{(t_{i})}$
$\mathcal{F}_{\theta}$ Multimodal fusion network
$f_{mm}^{(t_{\text{sys}})}$ Multimodal fused representation at system time $t_{\text{sys}}$
$\mathcal{H}_{\theta}$ Perception head (task-specific prediction layer)
$\delta_{i}^{\text{syn}}$ Clock drift (nominal temporal misalignment) for sensor $i$
$\delta_{i}^{\text{com}}$ Communication latency (nominal) for sensor $i$
$\delta_{i}^{\text{mal}}$ Added malicious delay for sensor $i$
$k$ Delay control threshold or parameter
${\mathcal{Y}}^{(t_{\text{sys}})}$ Predicted output (benign) at time $t_{\text{sys}}$
$\tilde{\mathcal{Y}}^{(t_{\text{sys}})}$ Predicted output (under attack or corrupted) at $t_{\text{sys}}$
$E_{mm}$ Shared encoder for aligned multimodal representations
$r_{i}^{(t_{i})}$ Projected shared representation of $x_{i}^{(t_{i})}$ from $E_{mm}$
$\lambda_{ij}$ Contrastive penalty weight between modalities $i$ and $j$
$\tau$ Temperature parameter for contrastive similarity scaling
$w$ Time window size used in detection or alignment
$\mathbf{S}$ Similarity matrix between representations
$\mathcal{P}^{*}$ Optimal temporal alignment path
$\phi^{*}$ Total alignment reward or cost along $\mathcal{P}^{*}$
$a$ Anomaly score from detection module
$n_{i}$ Index of the sample from $S_{i}$ at time $t_{i}$

### v2

@@
-On the Fragility of Multimodal Perception to Temporal Misalignment in Autonomous Driving
+Temporal Misalignment Attacks against Multimodal Perception in Autonomous Driving
 Abstract
@@
 However, its strict reliance on precise temporal synchronization exposes it to new vulnerabilities.
-In this paper, we introduce DejaVu, a novel attack that exploits network-induced delays to create subtle temporal misalignments across sensor streams, severely degrading downstream MMF-based perception tasks.
-Our comprehensive attack analysis across different models and datasets reveals these sensors’ task-specific imbalanced sensitivities: object detection is overly dependent on LiDAR inputs, while object tracking is highly reliant on the camera inputs.
+In this paper, we introduce DejaVu, an attack that exploits the in-vehicular network and induces delays across sensor streams to create subtle temporal misalignments, severely degrading downstream MMF-based perception tasks.
+Our comprehensive attack analysis across different models and datasets reveals the sensors’ task-specific imbalanced sensitivities: object detection is overly dependent on LiDAR inputs, while object tracking is highly reliant on the camera inputs.
 Consequently, with a single-frame LiDAR delay, an attacker can reduce the car detection mAP by up to 88.5%, while with a three-frame camera delay, multiple object tracking accuracy (MOTA) for car drops by 73%.
-To detect such attacks, we propose Aion, a defense patch that can work alongside the existing perception model to monitor temporal alignment through cross-modal temporal consistency.
-Aion leverages multimodal shared representation learning and dynamic time warping to determine the path of temporal alignment and calculate anomaly scores based on the alignment.
-Our thorough evaluation of Aion shows it achieves AUROC scores of 0.92–0.98 with low false positives across datasets and model architectures, demonstrating it as a robust and generalized defense against the temporal misalignment attacks.
+We further demonstrated two attack scenarios using an automotive Ethernet testbed for hardware-in-the-loop validation and the Autoware stack for end-to-end AD simulation, demonstrating the feasibility of the DejaVu attack and its severe impact, such as collisions and phantom braking.
 I Introduction
 Autonomous driving (AD) is designed to navigate and interact with complex environments—a capability fundamentally reliant on a comprehensive understanding of its surroundings.
-The integration of heterogeneous sensors, such as cameras, LiDARs, and radar, that capture data from different modalities allows an accurate perception with enhanced accuracy [1].
+The adoption of heterogeneous sensors, such as camera, LiDAR, and radar, that capture data from different modalities allows an accurate perception with enhanced accuracy and robustness [1].
 Each individual modality has unique strengths; for example, cameras capture rich semantic details, LiDAR provides accurate depth measurements, and radar particularly excels in detecting speed, even in adverse weather conditions [2].
 However, these sensors also face inherent limitations, such as a camera’s sensitivity to lighting variations, LiDAR’s lack of texture information, and radar’s sparsity, which can compromise performance when used independently [3, 4, 5].
-On the other hand, multimodal fusion (MMF) refers to the process of integrating such heterogeneous unimodal sensor data to generate a single and comprehensive representation of the surroundings for a more accurate perception [6], ensuring efficient downstream tasks such as object detection, tracking, and segmentation.
-MMF not only compensates for individual sensor weaknesses but also enhances the robustness of the overall system—a critical assumption that underpins its effectiveness in AD applications [7].
-Although MMF has demonstrated considerable effectiveness, it remains a complex problem and a fundamental research challenge due to the intrinsic heterogeneity of sensing modalities, due to unique data formats, resolutions, and noise characteristics [8, 9].
-This challenge is further compounded by the critical requirement of temporal alignment during sensor fusion, which is complicated by varying sampling rates across sensors, stemming from underlying hardware limitations [10].
-In safety-critical cyber-physical systems like AD, it is obligatory for real-time perception, decision-making, and control.
-On the other hand, temporal misalignments in sensor data acquisition—such as unexpected delays—can lead to significant degradation in the perception tasks, manifesting as object misdetection, inaccurate localization, and hence, scene misinterpretation.
-These perceptual inaccuracies undermine the reliability and safety of AD, adversely affecting core functionalities such as longitudinal and lateral control, maneuver execution, safety-critical interventions, etc.
-[11].
-These AD functions demand a perfectly synchronized sensor network with low latency to make timely control decisions [12].
-In that regard, Automotive Ethernet (AE) has emerged as a key networking technology for in-vehicular networks, offering data rates from 10 Mbps to 10 Gbps [13].
-Moreover, equipped with Time-Sensitive Networking (TSN), which uses generalized Precision Time Protocol (gPTP) for precise clock synchronization, ensures deterministic and real-time communication with bounded latency, which is essential for sensor fusion and control in AD applications [14, 15].
+Multimodal fusion (MMF)—the process of integrating multiple unimodal sensor data into a single and comprehensive representation—compensates for such individual sensor weaknesses, ensures accurate and robust perception, and efficient downstream tasks in AD [6, 7].
+MMF remains a fundamental research challenge, primarily due to the heterogeneity across sensing modalities, including differences in data formats, spatial resolutions, and temporal characteristics [8, 9].
+However, the robustness of temporal alignment—ensuring that all sensor data being fused corresponds to the same point in time—has received comparatively limited attention [10], despite being the fundamental enabler of MMF.
+Particularly, in AD, temporal alignment is a fundamental prerequisite to enable real-time perception, which directly informs decision-making and control.
+Misalignments, caused, for instance, by delays in specific sensor streams, can degrade perception performance, leading to object misdetection, localization errors, and scene misinterpretation.
+These inaccuracies can propagate downstream, affecting core functions such as control, maneuver planning, and safety interventions, ultimately compromising the reliability and safety of the vehicle [11].
+Moreover, temporal alignment in MMF for AD remains a complex problem with the following challenges.
 Figure 1: Impact of DejaVu attack on 3D object detection.
@@
 (c-d) Illustrate different temporal alignments between camera and LiDAR inputs and highlight how delayed sensor data can lead to incorrect detections, either by detecting non-existent objects (false positives) or missing present ones (false negatives).
-In (c), the MMF prioritizes the (delayed) LiDAR data and predicts three objects, including a pedestrian who is not present in the current camera view, resulting in a false detection from the camera’s perspective.
+In (c), the MMF prioritizes the (delayed) LiDAR data and predicts three objects, including a pedestrian who is not present in the current camera view, resulting in seemingly accurate results from LiDAR’s perspective but a false detection from the camera’s perspective.
 However, in (d), the MMF still prioritizes the (updated) LiDAR data and predicts two objects, excluding the pedestrian who is present in the current camera view, resulting in a missed detection from the camera’s perspective.
 In both temporal misalignment attack cases, MMF biased its fusion toward LiDAR, failing to account for semantic discrepancies in the camera modality.
-Although the adoption of AE ensures low latency and clock synchronization, temporal misalignment can still occur when data is captured at different sensors at different time points (timestamps), especially due to their varying sampling rates.
-Some industry-grade software stacks (such as Autoware [^1] https://github.com/autowarefoundation/autoware and Apollo [^2] https://github.com/ApolloAuto/apollo) process the unimodal data independently and use post-processing techniques, such as Kalman filters, to ensure temporal alignment during the fusion based on their corresponding timestamps.
-Therefore, the efficacy of MMF-based perception relies on two fundamental assumptions.
-First, $A_{1}$: AE with TSN is secure, which ensures the sensor networks are clock-synchronized and have low latency.
-Second, $A_{2}$: the timestamp attached to the packet is trustworthy and can be used during the fusion process.
-However, in this work we propose that an attacker can exploit at least any of $A_{1}$ or $A_{2}$ to compromise the sensor network and expose the MMF to temporal misalignment.
-MMF often relies on the assumption that integrating diverse sensor inputs will compensate for individual weaknesses and make it robust against any unimodal attack due to the redundancy of the collected data [8].
-However, research shows this redundancy does not inherently provide superior robustness.
-Both unimodal and multimodal systems can be vulnerable to carefully designed unimodal adversarial or patch attacks [16, 17, 18, 19, 20, 21].
-Moreover, research has shown that existing MMF-based perception models lack temporal robustness—the ability to produce reliable output despite temporal inconsistencies—against temporal misalignment and are quite fragile against a tiny delay, even in just one modality [10, 22, 23].
-Therefore, to overcome the high overhead of the adversarial or patch attacks but still jeopardize the perception of MMF, we propose a network-based temporal misalignment attack named DejaVu.
-Under the DejaVu attack, an attacker compromises the in-vehicle network and maliciously creates a misaligned temporal data stream to cause an illusion during perception, even without crafting data.
-The attacker can achieve this by breaking any of the assumptions that the reliability of MMF is based upon.
-First, the attacker can break $A_{1}$ by exploiting the existing vulnerabilities in gPTP, and desynchronize sensor networks.
-PTP has been shown to be vulnerable to various cyberattacks, including man-in-the-middle (MITM) attacks [24], packet delay attacks [25], and spoofing [26, 27].
-An attacker who gains such access to the in-vehicular network can manipulate PTP messages by introducing artificial delays, altering timestamps, or replaying old synchronization packets.
-This can cause different sensors to perceive capture data at slightly misaligned times, leading to misaligned sensor fusion.
-Similarly, the attacker can break $A_{2}$ by exploiting the known vulnerabilities in network infrastructure to introduce intensional delays or distortions in the timestamps, causing temporal misalignment in the pre-processing phase of MMF.
-Particularly, automotive networks typically use UDP [13], a protocol inherently susceptible to man-in-the-middle (MITM) attacks, where attackers can intercept and manipulate communication flows and timestamps without accessing the encrypted data to exacerbate the desynchronization [28, 29].
-In both cases, the attacker will be able to break the MMF assumptions of synchronized sensors or trustworthy timestamps, plummeting the perception accuracy.
+Clock synchronization.
+To ensure temporally aligned perception in AD, sensor data from all the modalities must be timestamped relative to a common global clock with minimal drift.
+In practice, the electronic controller units (ECUs) hosting the sensors must be tightly synchronized—often at sub-microsecond precision required by AUTOSAR [12])—to enable accurate fusion and downstream reasoning.
+To achieve that, automotive Ethernet (AE) equipped with Time-Sensitive Networking (TSN) has become the de facto backbone of modern in-vehicle networks.
+TSN leverages the generalized Precision Time Protocol (gPTP, IEEE 802.1AS) to maintain global clock synchronization among distributed ECUs [13, 14].
+Consequently, a core security assumption is: $\scriptstyle A_{1}$ The gPTP-based synchronization infrastructure is secure and maintains consistent global time across all ECUs.
+However, despite its widespread adoption in AE, gPTP was primarily designed to provide deterministic time synchronization—not to operate under adversarial conditions.
+Hence, gPTP lacks built-in security mechanisms, particularly cryptographic authentication, leaving it vulnerable to attacks such as grandmaster spoofing, delay injection, replay, and false time advertisement [15, 16, 17].
+Such attacks can compromise the temporal alignment of sensor fusion, undermining the integrity of MMF-based perception, without any direct alteration to the sensor data or sensor ECU itself.
+Timestamp Integrity.
+Even when sensors are nominally synchronized, they often capture data at slightly different times due to variations in sampling rates caused by inherent hardware limitations [10].
+Consequently, the timestamps attached to sensor messages become critical in multimodal fusion pipelines, as they serve as the primary reference for aligning asynchronous data streams by identifying the most temporally proximal pairs from buffered queues.
+This reliance gives rise to a fundamental security assumption: $\scriptstyle A_{2}$ Sensor ECUs are trusted entities that always provide accurate timestamps.
+However, this assumption can be violated.
+In vehicular systems, sensor ECUs may be compromised through remote exploits of unpatched software vulnerabilities [18, 19, 20], insecure Over-The-Air (OTA) updates [21, 22], or direct physical access via the OBD-II port [23, 24, 18].
+Once compromised, an attacker-controlled ECU can inject messages containing legitimate sensor data but with forged timestamps.
+Since the fusion ECU uses timestamps as the basis for temporal alignment, this can result in the selection of data pairs that appear temporally aligned but are actually misaligned, thereby degrading the integrity of the sensor fusion process.
+Middleware Integrity.
+Sensor data exchange in production-grade AD (such as Autoware [^1] https://github.com/autowarefoundation/autoware) is often facilitated by robotic middleware frameworks, such as robot operating system (ROS [^2] https://github.com/ros2/ros2), where the data distribution service (DDS) serves as the underlying communication backbone.
+DDS enables a real-time and scalable publish-subscribe communication model for data sharing among distributed ECUs, making it a widely adopted solution for managing the high-bandwidth, low-latency communication demands of multimodal perception and control pipelines.
+Consequently, a third foundational assumption arises: $\scriptstyle A_{3}$ The ROS-based DDS infrastructure is secure and ensures the integrity, authenticity, and freshness of shared data.
+However, the design of ROS prioritizes performance and scalability over robust security guarantees.
+Since ROS does not enforce strong authentication, encryption, or source verification by default, ROS is vulnerable to a wide range of network-level and application-level attacks, including message spoofing, replay, and impersonation attacks [25].
+As a result, an attacker with access to the ROS communication graph, for example, can impersonate legitimate nodes (e.g., a LiDAR publisher) and can publish fabricated sensor messages with targeted timestamps (i.e., replay attacks), which can be considered by the fusion ECU and mislead downstream perception tasks.
+Moreover, prior work has demonstrated that existing MMF-based perception systems in AD lack temporal robustness—the ability to maintain reliable outputs in the presence of temporal inconsistencies—making them vulnerable even to benign, non-malicious misalignments [10, 26, 27].
+In particular, perception pipelines have been shown to degrade significantly under small delays in just one modality.
+Existing studies, however, are limited in both scope and threat modeling: they primarily focus on random, system-induced delays and evaluate only a narrow subset of perception tasks—most commonly 3D object detection, without realizing them on any end-to-end AD software stacks.
+To investigate the true fragility of MMF-based perception under adversarial conditions, we present a comprehensive study of temporal misalignment attacks, which we term DejaVu.
+These attacks are realized by violating one or more of the core trust assumptions.
+Unlike prior work, we target both 3D object detection and multi-object tracking (MOT), and evaluate the effects of different delay distributions across multiple perception models and datasets.
 Fig.
 1 illustrates how temporal delays due to DejaVu attacks in one of the modalities can degrade the performance of the MVXNet model [28]—a MMF-based 3D object detection model, potentially leading to unsafe driving conditions in AD.
-The existing defense techniques against temporal misalignment assume benign timestamps and create alignment based on that.
-However, they will lack effectiveness against DejaVu, as the timestamps can be compromised as well.
-Hence, we propose a defense against DejaVu attack by extending this paradigm to include data content analysis along with the timestamps.
-In that regard, we propose Aion, a defense patch based on multimodal shared representation learning and dynamic time warping to detect and analyze the cross-modal temporal consistency.
-Aion works as a patch to any existing MMF models, and capitalizes on the semantic coherence of sensor data pairs, enabling the detection of inconsistencies that may arise from malicious temporal desynchronization.
 In summary, we make the following key contributions:
-  • We propose DejaVu, a novel temporal misalignment attack against MMF in AD that exploits vulnerabilities of in-vehicle networks and the fragility of multimodal fusion by selectively delaying sensor streams to disrupt perception in safety-critical tasks.
-  • We conduct a comprehensive empirical evaluation of DejaVu on state-of-the-art 3D object detection models (MVXNet [30], BEVFusion [31]) and a multi-object tracking model (MMF-JDT [32]) under various misalignment scenarios using the KITTI [33] and nuScenes [34] datasets.
+  • We propose DejaVu, a temporal misalignment attack against MMF in AD that exploits vulnerabilities of in-vehicle networks and the fragility of multimodal fusion by selectively delaying sensor streams to disrupt perception in safety-critical tasks.
+  • We conduct a comprehensive empirical evaluation of DejaVu on state-of-the-art 3D object detection models (MVXNet [28], BEVFusion [29]) using the KITTI [30] and nuScenes [31] datasets, and a multi-object tracking model (MMF-JDT [32]) under various misalignment scenarios using the KITTI [30] dataset.
 Our findings reveal distinct modality-specific vulnerabilities: object detectors are highly sensitive to LiDAR delays, while the tracking model is significantly impacted by camera timing disruptions.
 A single-frame LiDAR delay reduces 3D detection mAP by up to 88.5%, and a three-frame camera delay drops multiple object tracking accuracy (MOTA) by 73%.
-  • To mitigate these attacks, we introduce Aion, a defense patch that leverages multimodal shared representation learning and dynamic time warping to detect temporal inconsistencies.
-Aion achieves AUROC scores of 0.92–0.98 with a low false positive rate, demonstrating robust and effective detection of temporal misalignment.
+  • To further validate our findings in a realistic autonomous driving setting, we built an automotive Ethernet testbed that models the sensor data acquisition and fusion pipeline.
+Using this platform, we implemented the DejaVu attack by violating $\scriptstyle A_{2}$, demonstrating its feasibility in a hardware-in-the-loop environment.
+Additionally, to assess the end-to-end consequences beyond perception—specifically on planning and control—we integrated DejaVu into Autoware, a production-grade, full-stack autonomous driving simulator, by breaking $\scriptstyle A_{3}$.
+In both environments, our experiments show that DejaVu is highly practical and can result in severe safety violations, including direct collisions and phantom braking events.
+The remainder of the paper is organized as follows.
+Section II reviews related work; Section III outlines the threat model and proposes DejaVu attack; Section IV describes datasets and evaluation settings; Section V reports the evaluation results; Section VI reports the discussion and potential defenses; and Section VII concludes the paper.
 II Related Work
-II-A Attacks against MMF
-The attacks against MMF can be classified into the following two types:
-Sensor Spoofing Attacks.
-Sensor spoofing attacks focus on feeding false data to a vehicle’s sensors to mislead the autonomous system.
-For instance, LiDAR sensors can be fooled by malicious laser signals that fabricate or erase objects in the sensed environment [18].
-Beyond lasers, researchers have even created electromagnetic interference to tamper with LiDAR perception, achieving fake point insertions or deletions  [35].
-Camera sensors are likewise vulnerable; strategically placed adversarial patches on road signs can cause a ‘Stop sign’ to be ignored or misclassified as another sign [36].
-Such single-sensor spoofing attacks can trigger hazardous behaviors – e.g., a phantom obstacle can induce emergency braking, or a hidden obstacle can go unnoticed – underscoring how an attacker controlling one input modality can compromise the whole system.
-Such unimodal sensor spoofing attacks can be launched against multimodal model, while the remaining modalities remain unattacked.
-For example, the attacker can inject fake LiDAR points to create fake objects, aligning them to the camera’s view frustum [19], so it remains consistent with the camera image to avoid inconsistencies during the sensor fusion.
-On the other hand, it was assumed that an adversarial attack against both sensors at once is unlikely to be successful against AD perception that has multimodal inputs such as camera and LiDAR data.
-However, researchers have devised attacks that break this assumption by crafting inputs that confound multiple modalities simultaneously.
-It is possible to generate an adversarial 3D physical object that appears innocuous to humans but fools the fused camera+LiDAR detection pipeline [37].
-Attackers have even targeted high-level multi-modal representations by perturbing an image or audio input to make its fused embedding mimic an arbitrary target [38].
+The fundamental research direction has been in the direction of spoofing sensors from a single modality, such as LiDAR [33, 34, 35] and camera [36] through different means of physical perturbation.
+Moreover, advanced attacks have demonstrated to even compromise multiple modalities together [37, 38, 39].
 These sensor spoofing attacks demonstrate that simply having multiple sensors is not sufficient.
 With carefully constructed inputs, an adversary can simultaneously mislead camera and LiDAR sensors, defeating the very redundancy meant to ensure safety.
-Temporal Delay Attack.
-While time delay attacks have not been extensively studied within AD, they have created significant attention in other cyber-physical systems (CPS) domains, such as power systems, wireless networks, unmanned aerial vehicles (UAVs), robotics, and time-sensitive networks (TSNs).
-In power systems, the PTP is crucial for synchronization measurements and events.
-However, its vulnerability to time delay attacks poses risks to grid stability [39, 40].
-Wireless networks, particularly those used in industrial control systems, are susceptible to time delay attacks that can disrupt communication and control processes [41].
-In the realm of UAVs, time delay attacks can significantly impact flight control and mission execution [42, 43].
-Moreover, multimodal temporal misalignment in sensor fusion has been shown to degrade the accuracy of simultaneous localization and mapping (SLAM) [44].
-Software timing interference can also be exploited to cause system destabilization in CPS [45].
-Time-delay attacks also impact the traffic transmission quality under different combinations of Time-Aware Shaper (TAS) or Cyclic Queue Forwarding (CQF) in TSN [46].
-Furthermore, vulnerabilities in AE can allow for command injection attacks, leading to replay feed in camera [47].
-However, none of the existing works consider the impact of temporal misalignment against task-agnostic MMF-based perception.
+Unlike sensor spoofing attacks, time delay attacks have not been extensively studied within AD.
+They have created significant attention in other cyber-physical systems (CPS) domains, such as power systems [40, 41], wireless networks [42], unmanned aerial vehicles (UAVs) [43, 44, 15], and time-sensitive networks [45].
+Software timing interference is also exploited to cause system destabilization in CPS [46].
+Moreover, multimodal temporal misalignment in sensor fusion has been shown to degrade the accuracy of simultaneous localization and mapping (SLAM) [47], which was limited only to the fusion between IMU and camera data.
+Contrary to the existing works, we comprehensively study the impact of a temporal misalignment attack on task-agnostic MMF-based perception.
 Figure 2: Overview of the proposed system modeling with DejaVu attack.
-II-B Defense Strategies
-In the realm of autonomous vehicle security, various defense mechanisms have been proposed to counteract sensor spoofing and multimodal fusion attacks.
+In the realm of AD security, various defense mechanisms have been proposed to counteract sensor spoofing [48] and multimodal fusion attacks.
 These defenses can be broadly categorized into spatiotemporal consistency checks, specification-aware recovery strategies, and hardware-based techniques [49].
-One prominent approach involves detecting inconsistencies across different sensor modalities over time.
-In this defense mechanism, spatiotemporal consistency is checked to identify anomalies in sensor data, such as discrepancies in object motion or position between LiDAR and camera inputs [49].
-On the hardware side, manufacturers are improving sensor resilience: newer LiDAR models employ techniques such as pulse signal authentication or unique scanning patterns to make spoofed echoes easier to spot [20].
-However, none of these provide the means to detect temporal misalignment attacks in multimodal fusion for AD.
+PercepGuard [50] detects misclassification attacks by enforcing consistency between object tracks and class labels, but it does not examine the temporal validity of sensor readings and thus cannot detect replayed LiDAR scans whose semantic trajectories remain plausible.
+Connecting the Dots [51] employs class‐specific autoencoders to uncover context violations introduced by adversarial perturbations, yet time‐shifted data aligns perfectly with learned scene co‐occurrence statistics and evades its checks.
+PhyScout [52] formalizes cross‐modal conflict detection to identify gross spoofing, but subtle timestamp manipulations within the synchronizer’s tolerance window introduce no overt spatial or modality discrepancies.
+These approaches overlook timestamp integrity and data freshness as a security property.
 III Temporal Misalignment Attack : DejaVu
-This section elaborates on the system model, provides details of the proposed temporal misalignment attacks, and the threat model.
-Table II in the Appendix contains the summary of notations used throughout the paper.
-III-A System Modeling
-As shown in Fig.
-2, we consider an MMF-based perception system with $m$ sensors denoted as $S_{1},S_{2},\ldots,S_{m}$.
-At any time $t_{act}$, the transmitted sensor data from $S_{i}$ is represented by $(x_{i}^{(t_{act})},t_{pre})$.
-Definition 1 ($x_{i}^{(t_{\text{act}})},t_{\text{pre}}$).
-The tuple $(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$ represents a sensor data packet where $x_{i}^{(t_{\text{act}})}$ denotes the data “actually” captured at time $t_{\text{act}}$ (as indicated by the superscript), and $t_{\text{pre}}$ is the associated timestamp metadata, indicating the time at which the sample is “presumed” to have been captured.
-Discrepancies between $t_{\text{act}}$ and $t_{\text{pre}}$ suggest temporal misalignment.
-In MMF-based techniques timestamp $t_{pre}$ is used to ensure proper synchronization before fusing them and under a benign scenario at time $t_{i}$, the sensor data packet data can be expressed as $(x_{i}^{(t_{i})},t_{i})$, where $t_{i}=t_{act}=t_{pre}$.
-As each of the heterogeneous sensors generates data samples with a unique format and dimension, to ensure an effective sensor fusion, a modality-specific unimodal feature encoder (UFE) $E_{i}$ converts the raw data $x_{i}^{(t_{i})}$ to an intermediate representation $f_{i}^{(t_{i})}=E_{i}(x_{i}^{(t_{i})})$.
-Such encoders map the raw data to a similar feature format, which allows different sensor fusion techniques, such as straightforward operations like concatenation, merging, average polling, etc., or even more advanced techniques such as tensor-based fusion [50].
-Assume that sensor fusion is performed at a global time $t_{\text{sys}}$, using a fusion function denoted by $\mathcal{F}_{\theta}$.
-This process yields a fused representation $f_{mm}^{(t_{\text{sys}})}$, computed as follows:
-$f_{mm}^{(t_{\text{sys}})}=\mathcal{F}_{\theta}\big{(}f_{1}^{(t_{1})},f_{2}^{(t_{2})},\ldots,f_{m}^{(t_{m})}\big{)}$
-Finally, the corresponding perception head for the downstream tasks $\mathcal{H}_{\theta}$ predicts the perception results from the fused representation as:
-$\mathcal{Y}^{(t_{sys})})=\mathcal{H}_{\theta}\big{(}f_{mm}^{(t_{sys})}\big{)}$
-As $t_{i}$ is the actual time when the sensor $S_{i}$’s data was captured, and $t_{sys}$ is the time when the sensor data are synchronized and fused at the central ECU, under normal conditions, the $t_{i}$ can be expressed as:
-$t_{\text{sys}}=t_{i}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}}$ (1)
-where $\delta_{i}^{\text{syn}}$ represents the nominal temporal misalignment (clock drift) from the global time due to the clock drift of $S_{i}$.
-This delay is sensor-specific and under a perfect clock synchronization as mentioned in $A_{1}$; it becomes zero ($\delta_{i}^{\text{syn}}\approx 0$).
-On the other hand, $\delta_{i}^{\text{com}}$ is the communication latency, which indicates the time needed for the data to reach the sensor fusion module.
-With fast communication infrastructure (AE with TSN), this can be eliminated ($\delta_{i}^{\text{com}}\approx 0$).
-Therefore, under an ideal condition (such as $A_{1}$ $A_{2}$), $t_{sys}\approx t_{1}\approx t_{2}\approx\dots\approx t_{m}$ for $m$ sensors, ensuring that data from all sensors corresponds to the same system time $t_{sys}$.
-III-B Proposed Temporal Misalignment Attack: DejaVu
-In this part, we elaborate on our proposed temporal misalignment attack, named DejaVu attack, where an adversary maliciously introduces a timing delay $\delta_{i}^{\text{mal}}$ to one or more sensors with a goal of creating misaligned sensor fusion and creating a false perception of the surroundings.
-Under DejaVu attack, the attacker compromises the network to add an additional malicious delay of $\delta_{i}^{\text{mal}}$.
-Hence, the attacker transmits a compromised data packet $\big{(}\tilde{x}_{i}^{(\tilde{t}_{i})},t_{i}\big{)}$ to the fusion ECU with the outdated semantic content $\tilde{x}_{i}^{(\tilde{t}_{i})}$ which was captured at global time $\tilde{t}_{i}$ instead of $t_{i}$, but with the updated time stamp $t_{i}$.
-In this case, $t_{i}=\tilde{t}_{i}+\delta_{i}^{\text{mal}}$, and the attacker delayed that transmission by $\delta_{i}^{\text{mal}}$ either by breaking (such as $A_{1}$ $A_{2}$).
-Therefore, $t_{sys}$ can be expressed as:
-$t_{sys}=\underbrace{\tilde{t}_{i}+\delta_{i}^{\text{mal}}}_{\displaystyle{t_{i}}}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}}$ (2)
-The malicious delay $\delta_{i}^{\text{mal}}$ can vary over time–depending on the attacker’s intent, and be crafted to cause maximum disruption in sensor fusion.
-In a benign scenario, there will be no temporal misalignment, and sensor data will be received and processed in a timely manner.
-However, the attacker can launch different forms of DejaVu attack by controlling $\delta_{i}^{\text{mal}}$ in different ways.
-For instance, in Table I, we provide two possible DejaVu attacks scenarios against $S_{i}$, which are elaborated below:
-TABLE I: Two Types of DejaVu Attack Strategy Against $S_{i}$
-Attack Name Attack Type Delay Distribution $\delta_{i}^{\text{mal}}$
-Constant Delay Constant Constant, $k$
-Random Delay Random $\text{Uniform}(0,k)$
-Constant Delay
-This attack strategy introduces a fixed delay of $k$ frames in the sensor data stream.
-The attacker can achieve this by creating desynchronization among the clocks, tampering with sensor timestamps, or delaying data at the communication layer.
-Impact: The fusion model still receives temporally consistent data but with a lag in the target modality.
-This can lead to delayed (misaligned) perception (i.e., missing or shifted bounding boxes) in critical applications like object detection, where real-time perception is essential.
-Random Delay
-In this attack strategy, each frame experiences a different delay, randomly sampled from the range $[0,k]$.
-This strategy not only disrupts the real-time requirement but also disrupts the temporal sequence of sensor data, which is crucial for object tracking.
-Impact: The fusion system struggles to maintain the proper ordering of sensor inputs, leading to degraded perception accuracy.
-This will cause erratic behavior in time-sensitive sequential applications, particularly for object tracking and autonomous navigation.
-III-C Attackers Objective
-The adversary’s primary objective is to compromise the integrity and reliability of the sensor fusion by introducing deliberate temporal misalignment in one or more sensors.
-The attacker aims to disrupt the coherent integration of multi-modal sensor data, thereby inducing errors in perception.
-The resultant temporal inconsistencies can lead to incorrect perception results, particularly in safety-critical applications such as AD.
-Hence, the attacker’s key objective is to disrupt the perception to deteriorate the downstream tasks and hence, compromise the safety and integrity of the control decision.
-Based on the attacker’s capabilities, she can compromise either only one modality, which we define as unimodal DejaVu attack (Uni-DejaVu), or she can compromise multiple modalities together, defined as multi-modal DejaVu attack (Mul-DejaVu).
-Under the Uni-DejaVu attack with compromised, where only sensor $S_{1}$ is compromised, the compromised perception result $\tilde{\mathcal{Y}}^{(t_{sys})}$ at time $t_{sys}$ can be expressed as:
-$\tilde{\mathcal{Y}}^{(t_{sys})}=\mathcal{H}_{\theta}\big{(}\mathcal{F}_{\theta}\big{(}\tilde{f}_{1}^{(\tilde{t}_{1})},f_{2}^{({t}_{2})},\dots,f_{m}^{({t}_{m})}\big{)}\big{)}$ (3)
-Similarly, under the Mul-DejaVu attack with all the sensors compromised, the compromised perception result $\tilde{\mathcal{P}}(t_{sys})$ at time $t_{sys}$ can be expressed as:
-$\tilde{\mathcal{Y}}^{(t_{sys})}=\mathcal{H}_{\theta}\big{(}\mathcal{F}_{\theta}\big{(}\tilde{f}_{1}^{(\tilde{t}_{1})},\tilde{f}_{2}^{(\tilde{t}_{2})},\dots,\tilde{f}_{m}^{(\tilde{t}_{m})}\big{)}\big{)}$ (4)
-Under the DejaVu attack, the system will perceive the surroundings in a different way than that of the benign case $\tilde{\mathcal{Y}}^{(t_{sys})}\neq{\mathcal{Y}}^{(t_{sys})}$, which can lead to erroneous vehicle control decisions.
-III-D Attacker Capabilities
-We consider a sophisticated adversary capable of gaining access to the in-vehicle network, particularly AE, and launching targeted attacks that introduce temporal misalignment in MMF.
-We assume the attacker gains in-vehicle network access through one or more of the following realistic entry points [13]: Physical Access: The attacker connects to the vehicle’s OBD-II port or directly to Ethernet/CAN interfaces, either through malicious maintenance personnel or via physical compromise [51].
-Remote Exploitation: The attacker exploits vulnerabilities in externally exposed interfaces, such as telematics units, infotainment systems, or over-the-air (OTA) update mechanisms [51, 52].
-Supply Chain Attacks: The attacker implants malicious code or hardware during manufacturing, allowing persistent access post-deployment [53].
-Once access is established, the attacker can monitor, intercept, and inject messages on the in-vehicle Ethernet backbone and associated sub-networks.
-This enables manipulation of time-sensitive communications, particularly those involved in clock synchronization and timestamp propagation.
-We assume the attacker operates under two distinct capabilities.
-$C_{1}$ Disruption of Clock Synchronization.
-The adversary targets the clock synchronization mechanism in AE—specifically, the gPTP.
-Rather than altering timestamps directly, the attacker compromises the synchronization process itself, thereby inducing actual temporal misalignment between sensor streams.
-This can be accomplished by impersonating the grandmaster clock or tampering with synchronization messages via selective delay, replay, or man-in-the-middle (MITM) attacks [27, 24, 54].
-Although individual timestamps remain unaltered, they no longer correspond to a consistent global time due to induced drift.
-Consequently, sensor data arrives with nominally valid timestamps that are, in fact, temporally skewed, impairing the performance of the downstream perception task.
-$C_{2}$ Manipulation of Timestamp Integrity.
-In this scenario, the attacker is capable of preserving the actual timing of data transmission but alters the timestamps embedded in transmitted packets, particularly in UDP-based communication [13].
-This creates a seemingly temporal misalignment while data samples are, in fact, temporally aligned.
-As a result, while the fusion algorithm aligns the seemingly misaligned inputs, in reality, it creates actual temporal misalignment.
-Thus, we consider an attacker equipped with either $C_{1}$, $C_{2}$, or both capabilities to launch the DejaVu attack.
-IV Temporal Misalignment Defense: Aion
-Figure 3: Overview of the proposed defense Aion against any temporal misalignment attack.
-To defend against such temporal misalignment attacks, we propose a countermeasure technique named Aion, that can detect if any of the sensor data streams are misaligned.
-We design Aion as an independent detection patch that can work on top of any MMF-based application, either in parallel or sequentially, agnostic of the downstream task.
-Fig 3 shows the overview of Aion.
-In the following sections, we present an overview of the defense and each technical detail.
-IV-A Aion Overview
-As shown in Fig.
-3, Aion primarily consists of a single shared multimodal representation encoder (MRE) that maps any unimodal feature representation, regardless of its source or modality, to a shared representation space.
-Furthermore, Aion has two phases of implementation: i) development and ii) deployment phase.
-Development Phase.
-In the development phase, Aion trains an MRE model using contrastive learning with positive and negative pairs based on their related temporal alignment.
-MRE learns how to represent temporally aligned (positive) feature pairs from different modalities to similar representations and temporally misaligned (negative) pairs to different representations.
-Once the MRE is trained, Aion utilizes that trained MRE in the deployment phase to detect DejaVu attacks.
-Deployment Phase.
-During the deployment phase, Aion utilizes the trained MRE to create shared representations of historical inputs from each modality and keeps a stack of these representations for a small window.
-At the same time, Aion also calculates and keeps track of a 2D similarity matrix with pairwise inter-modality similarity scores between different representation pairs.
-The diagonal elements in the similarity matrix indicate pairs that are temporally aligned and others that are temporally misaligned to different extents as they deviate from the diagonal.
-On each such distance matrix, Aion runs a dynamic time warping (DTW) algorithm to find the optimal path of temporal alignment and the reward of such alignment, which is the summation of all their similarity scores.
-Under a benign scenario, the optimal path with the highest reward would be the diagonal one, and the reward would be higher.
-However, under a temporal misalignment attack, the optimal path would deviate from the diagonal and follow the attacker’s misaligned pattern.
-In that case, the optimal reward would be lower, which essentially indicates the existence of an adversary.
-We elaborate on the details of each component of Aion in the following subsections.
-IV-B Technical Details of Aion
-To learn a unified representation for multimodal inputs, we use a shared MRE that projects modality-specific features $f_{i}$ and $f_{j}$ from different modalities into a common latent space, such that $r_{i}^{(t_{i})}=E_{mm}(f_{i}^{(t_{i})})$ and $r_{j}^{(t_{j})}=E_{mm}(f_{j}^{(t_{j}})$.
-The objective is to ensure that the shared representations of semantically corresponding—i.e., temporally aligned—inputs are close in the latent space, meaning $r_{i}^{(t_{i})}=r_{j}^{(t_{i})}$ if $t_{i}=t_{j}$, and dissimilar otherwise.
-As the majority of MMF–based perception models for AD primarily focus on fusing camera and LiDAR data ($n=2$), we center our technical discussion of Aion on these two modalities.
-This choice aligns with common practice in the field.
-Thus, for the remainder of this section, we assume a two-modality setting involving sensors $S_{i}$ and $S_{j}$, which we refer to as $S_{C}$ and $S_{L}$, denoting camera and LiDAR, respectively.
-Nonetheless, the proposed Aion framework is designed to be modality-agnostic and readily generalizable to scenarios involving more than two sensor modalities.
-Moreover, under benign operating conditions, the sensors are assumed to be synchronized, with data from all modalities captured simultaneously at regular intervals.
-For notational convenience, instead of referring to the temporal sequences $t_{i}$ and $t_{j}$ which are continuous variables we consider their discrete time steps definied as $n_{i}$ and $n_{j}$, where they belong to the indices set $\mathcal{I}_{total}=\{1,2,3,\dots\}$.
-Thus, both $n_{i}$ and $n_{j}$ refer to the time indices for the corresponding sensors, with $r_{C}^{(n_{i})}$ indicating the shared representation of $(x_{C}^{(t_{i})},t_{i})$ , and $r_{L}^{(n_{j})}$ representing the shared representation of $(x_{L}^{(t_{j})},t_{j})$.
-IV-B1 Development Phase
-The development phase specifically involves the training of the MRE model and running the detection on benign data to set the threshold.
-To ensure effective learning, we utilize contrastive learning with three types of data pairs for the model training.
-Different Representation Pairs.  We categorize any representation pairs into three types based on their extent of temporal (mis-)alignment.
-Definition 2 (Positive Pairs).
-A pair $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})$ is considered a positive pair, denoted $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})\in\mathcal{T}_{p}$, if the corresponding sensor inputs are from the same temporal event, i.e., $n_{i}=n_{j}$.
-Definition 3 (Near-Negative Pairs).
-A pair $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})$ is considered a near-negative pair, denoted $(r_{C}^{(n_{i})},r_{L}^{(n_{j})}\in\mathcal{T}_{nn}$, if the corresponding sensor inputs are from different but temporally adjacent events, such that $n_{i}\neq n_{j}$ but $n_{i}\approx n_{j}$.
-In these pairs, $r_{C}^{(n_{i})}$ and $r_{L}^{(n_{j})}$ share partially overlapping semantic content due to their temporal proximity.
-Definition 4 (Far-Negative Pairs).
-A pair $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})$ is considered a far-negative pair, denoted $(r_{C}^{(n_{i})},r_{L}^{(n_{j})})\in\mathcal{T}_{fn}$, if the sensor inputs originate from totally distinct temporal events with no temporal or semantic overlap; that is, $n_{i}\neq n_{j}$ and $n_{i}\not\approx n_{j}$.
-Objective Function.
-The primary goal of the shared encoder $E_{mm}$ is to ensure that the representations of positive pairs are highly similar—i.e., have minimal distance—while representations of negative pairs remain well separated in the latent space.
-To achieve this, we adopt a contrastive learning objective, based on relaxed contrastive (ReCo) as proposed in [55], to train $E_{mm}$, where each training batch consists of a set of discrete sample indices $\mathcal{I}_{batch}=\{n_{1},n_{2},\dots,n_{b}\}$, where the batch size is $b$ and each $n_{k}$ corresponds to a unique sample in the batch.
-Thus, the representation sequences $\mathbf{r}_{C}=\{r_{C}^{(n_{1})},r_{C}^{(n_{2})},\dots,r_{C}^{(n_{b})}\}$ and $\mathbf{r}_{L}=\{r_{L}^{(n_{1})},r_{L}^{(n_{2})},\dots,r_{L}^{(n_{b})}\}$ from two different modalities are calculated on the sampled inputs from the training set.
-These indices are chosen in a manner that ensures the batch contains positive, near-negative, and far-negative pairs.
-Based on the $\mathbf{r}_{C}$ and $\mathbf{r}_{L}$, we compute a similarity matrix $\mathbf{S}\in\mathbb{R}^{b\times b}$, where each entry $S_{ij}$ denotes the cosine similarity between the camera representation $r_{C}^{(n_{i})}$ and the LiDAR representation $r_{L}^{(n_{j})}$, defined as:
-$S_{ij}=\frac{r_{C}^{(n_{i})}\cdot r_{L}^{(n_{j})}}{\|r_{C}^{(n_{i})}\|\,\|r_{L}^{(n_{j})}\|}$ (5)
-For positive pairs—i.e., pairs with maximum semantic alignment corresponding to the same temporal event—we expect the cosine similarity to approach 1.
-To enforce such alignment, we define the positive loss as:
-$\mathcal{L}_{\text{pos}}=\sum_{i=1}^{b}\left(S_{ii}-1\right)^{2}$ (6)
-This loss encourages the cosine similarity between the shared representations of temporally aligned inputs—i.e., $r_{C}^{(n_{i})}$ and $r_{L}^{(n_{i})}$—to be as close as possible to 1.
-Negative pairs consist of temporally misaligned inputs—either near-negatives (slightly misaligned) or far-negatives (strongly misaligned).
-Ideally, their representations should exhibit minimal cosine similarity.
-To enforce this, we define the negative loss as:
-$\mathcal{L}_{\text{neg}}=\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{b}\left(\max(0,S_{ij})\right)^{2}\cdot\lambda_{ij}$ (7)
-This loss penalizes cases where misaligned representations $r_{C}^{(n_{i})}$ and $r_{L}^{(n_{j})}$ exhibit undue similarity.
-The penalty is modulated by the weight $\lambda_{ij}$, which reflects the expected degree of dissimilarity based on temporal distance.
-For example, we hypothesize the weight $\lambda_{ij}$ based on the type of pair:
-$\lambda_{ij}=\begin{cases}\lambda\in(0,1),&\text{if }(r_{C},r_{L})\in\mathcal{T}_{nn}\quad\text{(near-negatives)}\\ 1,&\text{if }(r_{C},r_{L})\in\mathcal{T}_{fn}\quad\text{(far-negatives)}\end{cases}$
-To generalize this weighting scheme, we define $\lambda_{ij}$ as a smooth function of temporal distance:
-$\lambda_{ij}=\tanh\left(\frac{|n_{i}-n_{j}|}{\tau}\right)$ (8)
-Here, $\tau$ is a temperature-like scaling factor that controls sensitivity to temporal separation.
-This formulation offers a continuous and differentiable measure of misalignment, encouraging the model to learn nuanced distinctions across the temporal spectrum.
-Intuitively, near-negative pairs may retain some semantic overlap, and are thus penalized less for being similar.
-In contrast, far-negative pairs—expected to be semantically unrelated—are penalized more heavily if their representations remain close.
-This enables the model to learn robust, temporally-aware embeddings while maintaining sensitivity to semantic drift.
-The overall objective combines the positive and negative pair losses:
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{neg}}$ (9)
-This formulation promotes high cosine similarity for aligned (positive) pairs, while pushing apart misaligned (negative) pairs.
-The extent of separation for negative pairs is controlled by the penalty weight $\lambda_{ij}$, allowing for flexibility based on temporal misalignment.
-Algorithm 1 Optimal Warping Path and Reward Computation
-1:  Input: Cost matrix $\mathbf{S}\in\mathbb{R}^{w\times w}$
-2:  Output: Optimal warping path $\mathcal{P}^{*}$ and total optimal reward $R(w,w)$
-3:  Initialize accumulated similarity matrix $R\in\mathbb{R}^{w\times w}$
-4:  Set $R(1,1)=S(1,1)$Initialize first element cost}
-5:  for $n=2$ to $w$ do
-6:     $R(n,1)=S(n,1)+R(n-1,1)$
-7:  end for
-8:  for $m=2$ to $w$ do
-9:     $R(1,m)=S(1,m)+R(1,m-1)$
-10:  end for
-11:  for $n=2$ to $w$ do
-12:     for $m=2$ to $w$ do
-13:        $R(n,m)=S(n,m)+\max\left(R(n-1,m-1),R(n-1,m),R(n,m-1)\right)$
-14:     end for
-15:  end for
-16:  Backtracking to find the optimal warping path:
-17:  Initialize $\mathcal{P}^{*}=[(w,w)]$Start from the last position}
-18:  Set $(n,m)=(w,w)$
-19:  while $(n,m)\neq(1,1)$ do
-20:     if $n=1$ then
-21:        $m\leftarrow m-1$
-22:     else if $m=1$ then
-23:        $n\leftarrow n-1$
-24:     else
-25:        Choose $(n,m)$ from $\arg\max\left(R(n-1,m-1),R(n-1,m),R(n,m-1)\right)$
-26:     end if
-27:     Prepend $(n,m)$ to $\mathcal{P}^{*}$
-28:  end while
-29:  reward $\phi^{*}=R(w,w)$
-30:  Return: $\mathcal{P}^{*}$ and $\phi^{*}$
-IV-C Attack Detection
-The detection of DejaVu attacks, though Aion consists of two main tasks.
-First, Aion maintains individual queues of historical representations for each modality, and a similarity matrix indicating pair-wise similarity.
-Second, Aion analyzes the similarity matrix using dynamic time warping to find the optimal path of temporal alignment.
-From the observed optimal path and its total reward, Aion calculates the anomaly scores and detects any potential temporal misalignment.
-We elaborate on each step in the following subsections.
-IV-C1 Historical Representation Queue and Similarity Matrix
-Aion keeps queues of historical representations for each modality for the latest $w$ sensor data.
-If we assume the indices of that queue as $\mathcal{I}_{detect}=\{n_{1},n_{2},\dots,n_{w}\}$.
-For a presented window size $w$, Aion keeps track of the sequential representation $\mathbf{r}_{C}=\{r_{C}^{(n_{1})},r_{C}^{(n_{2})},\dots,r_{C}^{(n_{w})}\}$ and $\mathbf{r}_{L}=\{r_{L}^{(n_{1})},r_{L}^{(n_{2})},\dots,r_{L}^{(n_{w})}\}$.
-Similar to the training phase, as mentioned in (5), Aion creates the similarity matrix $\mathbf{S}\in\mathbb{R}^{w\times w}$.
-With the arrival of every new message, Aion updates $\mathbf{r}_{C}$, $\mathbf{r}_{L}$, and $\mathbf{S}$, and runs the DTW-based detection as described in the following part.
-IV-C2 Dynamic Time Warping-based Detection
-To quantify the extent of temporal misalignment within the $\mathbf{r}_{C}$ and $\mathbf{r}_{L}$, Aion employs DTW to compute both the optimal temporal alignment path and the corresponding alignment reward.
-We implemented DTW to identify the optimal warping path $\mathcal{P}$ that maximizes the accumulated similarity, which we define as reward, over a similarity matrix.
-Algorithm 1 outlines this procedure, which takes $\mathbf{S}$ as input and returns the optimal path $\mathcal{P}$ and total reward $\phi$ associated with that path.
-In an ideal scenario, where all sensors remain temporally aligned, the optimal warping path follows the diagonal: $\mathcal{P}^{*}=\{(1,1),(2,2),\dots,(w,w)\}$, as diagonal elements $S_{ii}$ have the highest similarity scores.
-Under the optimal alignment path, the optimal accumulated reward, $\phi^{*}=\sum_{i=1}^{w}S_{ii}=w$, since the embedding function $E_{mm}$ is trained to maximize similarity for aligned pairs.
-Thus, any deviation from that diagonal path $\mathcal{P}^{*}$ or the optimal reward $\phi^{*}$ can be considered anomalous.
-For any test case with optimal path $\mathcal{P}$ and total reward $\phi$, the anomaly score $a$ can be considered as $a=\phi^{*}-\phi=w-\phi$.
-In a benign scenario, all the sensors remain temporally aligned, and the optimal warping path usually follows the diagonal line ($\mathcal{P}_{ben}\approx\mathcal{P}^{*}$), as diagonal elements $S_{ii}$ have the highest similarity scores.
-Since $E_{mm}$ is trained to maximize similarity for aligned pairs, resulting in a high alignment reward $\phi_{ben}\approx w$ and low anomaly score $a_{ben}\approx 0$.
-Conversely, in the presence of a malicious actor in the network triggering temporal misalignment, the optimal warping path $\mathcal{P}_{mal}$ deviates from $\mathcal{P}^{*}$, incorporating non-diagonal elements where $n_{i}\neq n_{j}$, lowering the total reward $\phi_{mal}$ and hence, increasing the anomaly score $a_{mal}$.
-Justification on Detection.
-The fundamental assumption behind this approach is that DTW maximizes cumulative alignment reward by optimally aligning sequences.
-Given a well-trained $E_{mm}$, the cost function $S_{ij}$ satisfies:
-$S_{ij}=1\quad\text{iff}\quad n_{i}=n_{j}$
-In a benign case, where data samples are perfectly aligned, $\phi_{ben}$ is maximized, and $a_{ben}$ is minimized, since all elements on the optimal path mostly satisfy $n_{i}=n_{j}$, therefore:
-$\phi_{ben}=\sum_{(i,j)\in\mathcal{P}_{ben}}S_{ij}\approx\sum_{i=1}^{W}S_{ii}\quad\text{thus,}\quad a_{ben}\approx 0$
-However, in the presence of malicious misalignment, the warping path necessarily includes terms where $n_{i}\neq n_{j}$, leading to $S_{ij}<<1$ for some $(i,j)$.
-Since DTW optimally maximizes the total reward, the deviation from $\mathcal{P}^{*}$ implies a decrease in $\phi_{mal}$ and an increase in $a_{mal}$ is minimized, such that:
-$\phi_{mal}=\sum_{(i,j)\in\mathcal{P}_{mal}}S_{ij}<\sum_{i=1}^{W}S_{ii}\quad\text{thus,}\quad a_{mal}>>0$
-This establishes the fundamental assumption that as misalignment increases, so does anomaly score, reinforcing the validity of DTW in the anomaly detection process.
-Empirical validation in Section IV-C further supports this claim.
-V Evaluation on DejaVu
-We present the evaluation results of the DejaVu attack applied to MMF in the context of AD. This section details the experimental setup and discusses the key findings from the evaluation.
-V-A Experimental Settings
-We describe the experimental setup used to assess the DejaVu attack across different MMF models and datasets.
-V-A1 Datasets
-We experiment with two widely used multimodal autonomous driving datasets for both 3D object detection and multi-object tracking.
-KITTI Tracking Dataset.
-The KITTI Tracking Dataset [33] is a widely used benchmark for object detection and multi-object tracking in AD.
-It provides RGB images, LiDAR point clouds for a variety of urban, suburban, and highway driving scenarios for multiple object categories, including cars, pedestrians, and cyclists.
-NuScenes Dataset.
-The NuScenes Dataset [34] is a large-scale autonomous driving dataset that includes multi-modal sensor data, including six RGB cameras, a LiDAR with multiple categories, such as vehicles, pedestrians, and barriers.
-V-A2 Models
-This part elaborates on the state-of-the-art MMF models we consider to test the DejaVu attacks.
-To systematically evaluate the effect of DejaVu on MMF with different downstream tasks, we consider the following models:
-3D Object Detection: We evaluate DejaVu on two representative MMF-based 3D object detection models.
-i) MVXNet [30], trained on KITTI dataset, is an early fusion-based architecture that projects LiDAR point clouds into pseudo-image space and fuses them with camera image features at the voxel level.
-ii) BEVFusion [31] is a recent architecture, trained on the NuScenes dataset, unifies multi-modal sensor inputs—including LiDAR and camera—in the bird’s eye view (BEV) representation space.
-It employs modality-specific encoders followed by a fusion module that generates a shared BEV feature map for downstream detection.
-Multi-Object Tracking (MOT): For evaluating tracking performance under DejaVu, we consider MMF-JDT [32], trained on the KITTI tracking dataset, is a joint detection and tracking model that incorporates early and mid-level fusion strategies to align image and point cloud features for improved object association over time.
-V-A3 Attack Settings
-To evaluate the impact of DejaVu attack on MMF models, we introduce controlled delays in one or both sensor modalities (camera and LiDAR), as mentioned in Table I, and analyze the corresponding degradation in model performance.
-We systematically assess how different degrees and types of temporal misalignment affect the 3D object detection and multi-object tracking.
-We introduce both constant and random delays in one or both modalities, as defined in Table I.
-Specifically, we use a delay parameter $k\in\{0,1,2,3,4,5\}$, where $k=0$ represents a perfectly synchronized sensor, and $k=5$ denotes a maximum delay of five frames for the affected modality.
-We consider both Uni-DejaVu and Mul-DejaVu attacks.
-In Uni-DejaVu attack, delay is introduced in either the camera or the LiDAR input while keeping the other modality synchronized.
-In Mul-DejaVu attacks, both camera and LiDAR streams are delayed independently, leading to varying degrees of temporal misalignment.
-We use the pretrained weights for the target model provided with the official implementations.
-We focus exclusively on attacking the test dataset by applying the defined temporal delays.
-V-A4 Evaluation Metrics
-To assess the impact of DejaVu, we analyze the model performance using task-specific evaluation metrics.
-For 3D object detection, we evaluate the models using mean Average Precision (mAP), which quantifies the accuracy of detected objects.
-For MOT, we use standard tracking metrics such as higher order tracking accuracy (HOTA), detection accuracy (DetA), association accuracy (AssA), multiple object tracking accuracy (MOTA), and Identity Switches (IDSW), which measure the effectiveness of object association across frames.
-V-B Software Implementation
-We implement both DejaVu and Aion using Python 3.8 and PyTorch, utilizing open-source frameworks including MMDetection3D [56] and OpenPCDet [57].
-Experiments were conducted on a server running Ubuntu 20.04.6 LTS with an Intel Xeon Gold 5520 (16 cores, 2.20GHz), 128GB RAM, and three NVIDIA RTX 6000 Ada GPUs.
-V-C Evaluation Results
-This section presents the DejaVu attack impact on different MMF models and datasets.
-V-C1 Impact of DejaVu Attack on 3D Object Detection
-We investigate the effectiveness of the proposed DejaVu attack on multimodal 3D object detection using MVXNet and BEVFusion, evaluated on the KITTI and nuScenes datasets, respectively.
-We analyze the detection performance of 3D object detection across different object classes under varying levels of unimodal and multimodal sensor delay.
-Object detection models do not process sequential information; instead, their performance is affected only by frame-wise delays in each modality at any particular time, whether the delays are constant or random.
-Therefore, for simplicity and consistency, we only analyze the attack under the constant delay setting.
-MVXNet on KITTI Dataset
-Fig 4 presents the 3D object detection performance of MVXNet under Uni-DejaVu attacks, where either the camera or LiDAR input is delayed independently.
-Under benign (zero-delay) conditions, MVXNet achieves strong mAP across most object classes: approximately 84.1 for cars, 87.6 for pedestrians, and 73.2 for cyclists.
-The left plot shows that delaying the camera input alone–under Uni-DejaVu camera attacks—has minimal effect on performance across all classes, with nearly constant mAPs.
-In contrast, the right plot highlights the model’s high sensitivity to LiDAR delays: a 1-frame delay causes the car mAP to collapse from 84.1 to 9.7 ($\downarrow$88.5%), pedestrian mAP from 87.6 to 34.2 ($\downarrow$60.9%), and cyclist mAP from 73.2 to 32.9 ($\downarrow$55.1%), with further degradation as delay increases.
-This indicates that LiDAR data is significantly more critical than camera input in MVXNet’s perception pipeline; hence, MVXNet’s high vulnerability against Uni-DejaVu attack against LiDAR.
-However, Fig.
-5 shows the mAP heatmaps across combinations of camera and LiDAR delays under Mul-DejaVu attacks.
-Although the 1-frame LiDAR delay drops mAP from 55.1% to 88.5% for different objects, further camera delay along with LiDAR has no additional effect, indicating the dominance of LiDAR in MVXNet.
-Key Findings: MVXNet heavily depends on LiDAR input for 3D object detection.
-Camera delay, under both Uni- or Mul-DejaVu attack, has almost no effect, but even minor LiDAR misalignment leads to severe performance degradation.
-When both modalities are delayed, the impact is still only dominated by the LiDAR stream.
-Figure 4: Uni-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
-Figure 5: Mul-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
-Figure 6: Uni-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
-Figure 7: Mul-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
-(a) Constant Delay Scenario
-(b) Random Delay Scenario
-Figure 8: DejaVu attack impacts on multi-object (car) tracking performance of MMF-JDT on KITTI tracking dataset with respect to different metrics.
-BEVFusion on nuScenes Dataset
-Fig.
-6 shows BEVFusion’s 3D detection performance under Uni-DejaVu attacks.
-With no delay, the model achieves high mAP across most classes—approximately 88 for car and pedestrian, and slightly lower for the rest.
-With a 1-frame camera delay, mAP drops slightly across all classes (for instance, car mAP drops by 7.4%), but remains stable with further delays, showing BEVFusion is slightly vulnerable against Uni-DejaVu camera attacks.
-In contrast, introducing a 1-frame LiDAR delay results in a substantial reduction in mAP for specific object classes, with performance dropping by 65.6% for cars (from 88.8 to 30.5), 35.1% for pedestrians (from 88.1 to 57.2), and 38.6% for motorcycles (from 75.9 to 46.6).
-Although performance remains stable with further LiDAR delays, this finding shows BEVFusion’s higher vulnerability against Uni-DejaVu LiDAR attacks.
-Similarly, Fig.
-7 illustrates BEVFusion’s performance under Mul-DejaVu attacks.
-Although delaying the camera alone has a minimal impact on performance, combining this with a delay in the LiDAR stream causes a significant drop in mAP.
-For instance, a 1-frame delay in camera or LiDAR stream reduces car mAP by 7.4% and 65.5%, respectively.
-However, when both modalities are delayed simultaneously by one frame, the mAP drops dramatically by 89.2% (from 88.8 to 9.6), highlighting a substantial compounding effect.
-This trend is consistently observed across all object classes.
-Key Findings.
-BEVFusion is slightly affected by camera delays but is highly affected by LiDAR delays, further underscoring LiDAR’s dominating role in 3D object detection.
-However, the impact becomes significant if both sensors are delayed simultaneously, even just by one frame.
-V-C2 Impact of DejaVu Attack on Multi Object Tracking
-We investigate the effectiveness of the proposed DejaVu attack on MOT algorithm using MMF-JDT on the KITTI dataset.
-We analyze the tracking performance for cars under varying levels of delays under Uni- and Mul-DejaVu attack scenarios.
-MMF-JDT on KITTI Dataset
-Fig 8 illustrates the impact of DejaVu attacks on MMF-JDT evaluated on the KITTI tracking dataset.
-For this evaluation, we consider both types of delays: constant (Fig 8a) and random (Fig 8b).
-Across both attack scenarios, the tracking performance declines as camera delay increases.
-Along with other tracking metrics, MOTA and IDSW suffer noticeable degradation as camera delays increase under Uni-DejaVu attacks.
-For instance, IDSW—a metric that captures identity switches and ideally should be low, increases dramatically under increasing camera delays, underscoring the disruption in tracking consistency caused by DejaVu attacks.
-Furthermore, the values of different metrics across the heatmaps of Fig 8a suggest that while camera delay under Uni-DejaVu deteriorates the performance, delaying both the camera and LiDAR by exactly the same delay (i.e., constant delay scenario) under Mul-DejaVu attack diminishes the attack impact and mostly retains the performance.
-On the other hand, heatmaps in Fig.
-8b show that Mul-DejaVu attacks with random delays remain effective as different delays in both modalities break the sequence, making object tracking considerably more difficult.
-For instance, under the Mul-DejaVu constant attack scenario with a five-frame delay, MOTA decreases by only 1.4%, whereas the random attack scenario results in a 60.5% drop.
-Key Findings.
-In contrast to 3D object detection tasks, MOT appears to rely more heavily on camera inputs.
-This may be attributed to the fact that MOT does not require precise 3D bounding boxes; instead, the rich texture information in camera images may offer more effective contrastive representations than sparse point clouds.
-As a result, DejaVu attacks can substantially impair MOT performance, particularly under Uni-DejaVu (camera delay) and Mul-DejaVu with random delay scenarios.
-VI Evaluation of Aion
-VI-A Experimental Settings
-To evaluate the effectiveness of Aion in detecting DejaVu attacks, we conduct a detection analysis under various attack scenarios.
-We synthetically generate different degrees of temporal misalignment by perturbing the input sequences in the test data as described in Table I.
-For two different models trained on two different datasets, we evaluate Aion’s ability to distinguish between normal and misaligned sequences under diverse DejaVu attacks.
-Model Architecture of Aion
-We implemented Aion for both the KITTI and nuScenes datasets to evaluate its adaptability across different sensor setups and driving scenarios.
-Aion on KITTI: For the KITTI dataset, we adopt a straightforward approach by testing with two off-the-shelf, pre-trained image and LiDAR feature encoders.
-The MRE of Aion is implemented using a simple convolutional neural network (CNN) architecture, featuring two distinct input branches and a shared output branch.
-For the KITTI dataset, each input sample consists of an RGB image and a corresponding LiDAR point cloud.
-The RGB input image has the size of $[3,375,1242]$, and is processed using a pretrained ResNet-50 [58] feature encoder, yields image feature maps $f_{C}$ of shape $[2048,12,39]$.
-Simultaneously, the LiDAR input of size $[k,3]$, where $k$ is the number of points, is voxelized using a pretrained PointPillars [59] encoder, producing LiDAR feature maps $f_{L}$ of shape $[384,248,216]$.
-Later, we use our proposed $E_{mm}$ to project both $f_{C}$ and $f_{L}$ to a shared space with representations $r_{C}$ and $r_{L}$, respectively.
-First, both $f_{C}$ and $f_{L}$ are passed through their modality-specific convolutional input branch that reduces their channel dimension to 256, with progressive spatial downsampling, followed by a global average pooling layer, and a shared fully connected projection head to produce 256-dimensional representation vectors for both $r_{C}$ and $r_{L}$.
-Figure 9: Similarity scores between Camera and LiDAR representation embeddings under benign and DejaVu attacks between time steps 25 to 35.
-Aion on nuScenes: For the nuScenes dataset, we build Aion directly on top of BEVFusion [31] architecture to showcase Aion’s adaptability with a more complex MMF architecture.
-Due to the complexity of the nuScenes dataset, where each input contains six camera images and a 360 lidar point cloud, we introduce a hybrid architecture for $E_{mm}$, utilizing both CNN and transformer architectures, to learn compact and robust representations.
-Using BEVFusion’s own individual input feature encoders, we analyze the six images and create a unified BEV feature map $f_{C}$ with size $[64,180,180]$.
-Point clouds from lidar are also processed and mapped to the feature vector $f_{L}$ with the same dimension using the slightly modified point cloud encoder, as detailed in [31].
-Later, $f_{C}$ and $f_{L}$ are fed to our proposed hybrid $E_{mm}$, which has two parts.
-First, they are passed through a shared convolutional pipeline composed of three CNN layers to have an intermediate embedding with output shape $[256,23,23]$.
-Later, we adopt a transformer architecture to process the intermediate representation, where the feature maps are reshaped into spatial tokens with $256$-dimensional embeddings, and positional encodings are added.
-A lightweight transformer encoder with a single self-attention layer enables global spatial reasoning.
-The sequence is reduced using mean pooling, resulting in $r_{C}$ or $r_{L}$, which are $256$-dimensional shared representations.
-For both the KITTI and nuScenes datasets, we train the $E_{mm}$ with multimodal contrastive learning, as mentioned in Section IV-B1, and train the models for $100$ and $500$ epochs, respectively, on the training dataset.
-While training $E_{mm}$ we consider $\tau=3$, which sets the boundary of near- and far-negative sample pairs, during the loss calculation, as mentioned in (7) and (8).
-Figure 10: Visualization of anomaly scores generated by Aion under different DejaVu attacks on different modalities.
-The ‘red’ and ‘green’ shaded areas indicate the time periods with or without DejaVu attacks, respectively.
-The distinctive anomaly scores at different regions show Aion’s high effectiveness.
-Attack Synthesis
-For both datasets, we synthetically launch DejaVu attacks on the test data sequences at certain intervals to create malicious test sequences with varying degrees of misalignment.
-Specifically, we launch each attack scenario outlined in Table I at every $25$ time steps, which persist for the next $k=10$ continuous time steps.
-This systematic attack injection ensures a dataset containing misaligned samples with different extents of temporal shifts, allowing for a comprehensive evaluation of the defense mechanism.
-Anomaly Detection Methodology.
-To classify whether an input sequence is malicious, we analyze the cross-modal temporal consistency of multimodal pairs within a defined observation window $w=5$.
-In this evaluation, we label a window as malicious if at least one of its multimodal pairs contains a misaligned sample.
-We analyze the anomaly scores using the ROC curve and calculate the area under the ROC curve (AUROC) as the key detection metric.
-VI-B Detection Results
-We evaluate the performance of Aion across both datasets and model architectures.
-We begin by illustrating the detection process on the nuScenes dataset, including visualizations of similarity and anomaly scores under different attack types.
-Finally, we present the ROC curves, along with the AUROC scores, for both datasets.
-VI-B1 Visualization of Similarity Matrix
-Fig 9 illustrates four different similarity matrices with the Camera and LiDAR representations, $r_{C}$ and $r_{L}$, from time steps 10 to 50 under various DejaVu attacks (launched from 25 to 35), including the benign case.
-The left-most panel shows the similarity matrix between $r_{C}$ and $r_{L}$ under benign conditions—i.e., with no delay in either modality.
-As illustrated, the highest similarity scores lie along the diagonal path from (10, 10) to (50, 50), indicating perfect temporal alignment between both modalities.
-However, the two middle panels depict cases where two types of temporal misalignments are introduced under Uni-DejaVu attacks: one with a constant delay applied to the camera stream, and another with a random delay introduced in the LiDAR stream, both between time steps 25 and 35.
-In these scenarios, the highest similarity scores diverge from the diagonal beyond time step 25 and only return to the diagonal again around time step 35.
-These deviations clearly signify temporal misalignments, which Aion leverages to detect such DejaVu attacks.
-The right-most panel presents a unique scenario where both modalities are delayed by the same amount (constant delay) under Mul-DejaVu attack.
-In this case, the similarity scores remain high (and the same) across both diagonal and off-diagonal elements from time steps 25 to 35.
-Such patterns may emerge under both benign and malicious conditions.
-For instance, under benign conditions, the vehicle may be stationary without any moving objects in the scene, resulting in temporally consistent features over time.
-In contrast, an attacker could also replicate this same scene with malicious delay to all of the modalities by the same constant offset, creating a similar similarity matrix.
-Hence, these unique advanced attack becomes a challenging task just by analyzing the cross-modal alignment similarities.
-Although Aion, when limited to only the modalities used in MMF, cannot reliably detect such an advanced attack case, incorporating additional data sources—such as inertial measurements (IMU), controller area network (CAN) signals, or other external references—can provide complementary evidence and help detect such advanced attacks.
-However, as we only rely on the multimodal data in this work, we consider this extension as a future work for Aion.
-VI-B2 Detection Performance of Aion
-We illustrate the detection performance of Aion from two different perspectives.
-Visualization of Anomaly Scores
-Fig.
-10 illustrates the temporal evolution of anomaly scores, provided by Aion, across different time steps under various attack scenarios.
-Each shaded region indicates whether the system is operating under benign (green) or malicious (red) conditions, based on the temporal alignment.
-As shown, Aion consistently produces higher anomaly scores during periods where temporal misalignment is introduced, compared to benign intervals where no such misalignment is present.
-This clear contrast in anomaly demonstrates the effectiveness of Aion in detecting malicious temporal misalignment induced by DejaVu attacks on different modalities.
-ROC Curve with AUROC Scores
-Here, we provide quantitative evaluations of Aion’s detection performance on both nuScenes and KITTI datasets.
-Fig.
-11 presents the ROC curves evaluating the performance of the Aion in detecting DejaVu attacks on the nuScenes dataset under various sensor delay scenarios.
-The evaluation encompasses both constant and random delay injections on camera, lidar, and both modalities.
-Aion achieves consistently high AUROC scores, ranging from $0.94$ to $0.97$, demonstrating strong detection capabilities.
-In particular, constant delays on the camera and lidar sensors yield the highest AUC of $0.97$.
-Even under more challenging conditions, such as random delays affecting both sensors, the Aion maintains an AUC of $0.94$.
-These results illustrate that the Aion effectively detects temporal misalignment under DejaVu attacks while maintaining low false positive rates.
-Figure 11: ROC curve of the Aion against DejaVu attacks on nuScenes dataset.
-Figure 12: ROC curve of the Aion against DejaVu attacks on KITTI dataset.
-Fig.
-12 illustrates the ROC curves corresponding to the DejaVu detection mechanism on the KITTI dataset across identical attack scenarios as considered above.
-Similar to nuScenes, Aion also demonstrates high efficacy against KITTI dataset, achieving AUROC values up to $0.98$ for constant delay attacks on camera and lidar individually.
-In the presence of random delay attacks against one or both modalities, the detection performance remains robust, with AUROC scores ranging from $0.92$ to $0.97$.
-These findings substantiate the generalizability of the Aion across datasets and model architectures, further emphasizing its practical effectiveness.
-Moreover, the consistently high true positive rates with a low false positive rates underscore the Aion’s reliability in realistic AD environments subjected to DejaVu attacks.
-VII Conclusion
-This work presents DejaVu, a novel temporal misalignment attack that exploits synchronization vulnerabilities in multimodal perception systems for autonomous driving.
-Through extensive evaluations on state-of-the-art 3D object detection and multi-object tracking models, we uncover modality-specific vulnerabilities: 3D detection models are predominantly reliant on LiDAR and suffer severe degradation—with up to 88.5% drop in mAP—from even a single-frame LiDAR delay, while MOT models exhibit heightened sensitivity to camera stream disruptions, with MOTA dropping by 73% under just three-frame camera delays.
-To mitigate these vulnerabilities, we propose Aion, a detection framework based on multimodal shared representation learning and dynamic time warping.
-Aion consistently achieves AUROC scores of 0.92–0.98 with low false positive rates on both KITTI and nuScenes datasets.
-These findings highlight the critical need for synchronization-aware design in perception architectures and emphasize the importance of robust temporal consistency checks in safety-critical autonomous systems.
-Appendix A Overall Evaluation Results on BEVFusion
-Fig. 13 shows the overall impact of Mul-DejaVu on average mAP and NDS score on the nuScenes dataset.
-Figure 13: BEVFusion (nuScnenes): Overall 3D object detection mAP under Mul-DejaVu attacks for different object classes regarding i) NDS, ii) mAP
-Appendix B Notation Summary
-TABLE II: Summary of Notations
+This section elaborates on the system model, the threat model, and the details of the proposed temporal misalignment attacks. Table I contains a summary of notations used throughout the paper.
+TABLE I: Summary of Notations
 Symbol Description
@@
 $\tilde{\mathcal{Y}}^{(t_{\text{sys}})}$ Predicted output (under attack or corrupted) at $t_{\text{sys}}$
-$E_{mm}$ Shared encoder for aligned multimodal representations
-$r_{i}^{(t_{i})}$ Projected shared representation of $x_{i}^{(t_{i})}$ from $E_{mm}$
-$\lambda_{ij}$ Contrastive penalty weight between modalities $i$ and $j$
-$\tau$ Temperature parameter for contrastive similarity scaling
-$w$ Time window size used in detection or alignment
-$\mathbf{S}$ Similarity matrix between representations
-$\mathcal{P}^{*}$ Optimal temporal alignment path
-$\phi^{*}$ Total alignment reward or cost along $\mathcal{P}^{*}$
-$a$ Anomaly score from detection module
-$n_{i}$ Index of the sample from $S_{i}$ at time $t_{i}$
+III-A System Modeling
+As shown in Fig.
+2, we consider an MMF-based perception system with $m$ sensors denoted as $S_{1},S_{2},\ldots,S_{m}$.
+At any time $t_{\text{act}}$, the transmitted sensor data from $S_{i}$ is represented by $(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$.
+Definition 1 ($x_{i}^{(t_{\text{act}})},t_{\text{pre}}$).
+The tuple $(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$ represents a sensor data packet where $x_{i}^{(t_{\text{act}})}$ denotes the data “actually” captured at time $t_{\text{act}}$ (as indicated by the superscript), and $t_{\text{pre}}$ is the associated timestamp metadata, indicating the time at which the sample is “presumed” to have been captured.
+Discrepancies between $t_{\text{act}}$ and $t_{\text{pre}}$ suggest temporal misalignment.
+In MMF-based techniques timestamp $t_{\text{pre}}$ is used to ensure proper synchronization before fusing them and under a benign scenario at time $t_{i}$, the sensor data packet data can be expressed as $(x_{i}^{(t_{i})},t_{i})$, where $t_{i}=t_{\text{act}}=t_{\text{pre}}$.
+As each of the heterogeneous sensors generates data samples with a unique format and dimension, to ensure an effective sensor fusion, a modality-specific unimodal feature encoder (UFE) $E_{i}$ converts the raw data $x_{i}^{(t_{i})}$ to an intermediate representation $f_{i}^{(t_{i})}=E_{i}(x_{i}^{(t_{i})})$.
+Such encoders map the raw data to a similar feature format, which allows different sensor fusion techniques, such as straightforward operations like concatenation, merging, average polling, etc., or even more advanced techniques such as tensor-based fusion [53].
+Assume that sensor fusion is performed at a global time $t_{\text{sys}}$, using a fusion function denoted by $\mathcal{F}_{\theta}$.
+This process yields a fused representation $f_{mm}^{(t_{\text{sys}})}$, computed as follows:
+$f_{mm}^{(t_{\text{sys}})}=\mathcal{F}_{\theta}\big(f_{1}^{(t_{1})},f_{2}^{(t_{2})},\ldots,f_{m}^{(t_{m})}\big)$
+Finally, the corresponding perception head for the downstream tasks $\mathcal{H}_{\theta}$ predicts the perception results from the fused representation as:
+$\mathcal{Y}^{(t_{\text{sys}})})=\mathcal{H}_{\theta}\big(f_{mm}^{(t_{\text{sys}})}\big)$
+As $t_{i}$ is the actual time when the sensor $S_{i}$’s data was captured, and $t_{\text{sys}}$ is the time when the sensor data are synchronized and fused at the central ECU, under normal conditions, the $t_{i}$ can be expressed as:
+$t_{\text{sys}}=t_{i}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}},\quad\forall i\in\{1,\dots,m\}$ (1)
+where $\delta_{i}^{\text{syn}}$ represents the nominal temporal misalignment (clock drift) from the global time due to the clock drift of $S_{i}$.
+This delay is sensor-specific and under a perfect clock synchronization, it becomes zero ($\delta_{i}^{\text{syn}}\approx 0$).
+On the other hand, $\delta_{i}^{\text{com}}$ is the communication latency, which indicates the time needed for the data to reach the sensor fusion module.
+With fast communication infrastructure (AE with TSN) with efficient middleware, this can be eliminated ($\delta_{i}^{\text{com}}\approx 0$).
+Therefore, under an ideal condition (such as $\scriptstyle A_{1}$$\scriptstyle A_{3}$), $t_{\text{sys}}\approx t_{1}\approx t_{2}\approx\dots\approx t_{m}$ for $m$ sensors, ensuring that data from all sensors corresponds to the same system time $t_{\text{sys}}$.
+III-B Threat Model
+III-B1 Attacker Capabilities
+In the threat model with an adversary capable of gaining access to the in-vehicle network through one or more of the following realistic entry points [54]: Physical Access: The attacker connects to the vehicle’s OBD-II port or directly to Ethernet/CAN interfaces, either through malicious maintenance personnel or via physical compromise [18].
+Remote Exploitation: The attacker exploits vulnerabilities in externally exposed interfaces, such as telematics units, infotainment systems, or over-the-air (OTA) update mechanisms [18, 20, 22].
+Supply Chain Attacks: The attacker implants malicious code or hardware during manufacturing, allowing persistent access post-deployment [55].
+Once access is established, the attacker can monitor, intercept, and inject messages on the in-vehicle AE backbone and associated sub-networks.
+This enables manipulation of time-sensitive communications, particularly those involved in clock synchronization and forged data/timestamp propagation.
+We assume the attacker operates under any of these three distinct capabilities.
+$\scriptstyle C_{1}$
+Disruption of Clock Synchronization.
+With this capability, the adversary targets the clock synchronization mechanism in AE—specifically, the gPTP.
+Rather than altering timestamps directly, the attacker compromises the synchronization process itself, thereby inducing actual temporal misalignment between sensor streams and other sub-networks.
+This can be accomplished by impersonating the grandmaster clock or tampering with synchronization messages via selective delay, replay, or man-in-the-middle attacks [56, 57, 58].
+Although individual timestamps remain unaltered, they no longer correspond to a consistent global time due to induced drift, impairing the performance of the downstream perception task.
+$\scriptstyle C_{2}$
+Manipulation of Timestamp Integrity.
+In this scenario, the attacker is capable of preserving the actual timing of data transmission but alters the timestamps embedded in transmitted packets [54].
+This creates a seemingly temporally misaligned messages in the data queue from the timestamps perspective while the data contents are, in fact, temporally aligned.
+Consequently, while the data synchronizer [^3] TimeSynchronizer and ApproximateTimeSynchronizer are commonly used message filtering utilities in ROS2 that align multiple sensor message streams based on their timestamps.
+While TimeSynchronizer performs strict timestamp matching, ApproximateTimeSynchronizer allows messages with slight temporal differences—within a specified tolerance window—to be synchronized.
+finds the seemingly aligned inputs, in reality, it finds actual temporal misaligned pairs.
+$\scriptstyle C_{3}$
+Impersonation of a Legitimate Node in ROS2.
+In this scenario, the attacker is assumed to be a participant in the ROS2 network.
+As the default ROS2 implementation lacks built-in security mechanisms—a configuration commonly adopted in industry-grade AD software stacks, including Autoware—the attacker can instantiate a malicious node that impersonates a legitimate sensor publisher.
+By subscribing to target sensing topics [^4] Representative topic names from Autoware are /sensing/camera/traffic_light/image_raw, /sensing/lidar/top/pointcloud_raw, etc., the attacker gains access to the data stream and records historical sensor messages.
+The malicious node then re-publishes previously captured messages with updated, genuine-looking timestamps, potentially while the original publisher is still active.
+By repeatedly injecting such delayed-but-valid messages, the attacker can pollute the input queue of time-based synchronizers, increasing the likelihood that a forged message is selected during the fusion process.
+Thus, we consider an attacker equipped with either of these capabilities or even multiple of them to launch the DejaVu attack, thereby causing temporal misalignment and compromising perception integrity.
+III-B2 Attackers Objective
+The adversary’s primary objective is to compromise the integrity and reliability of the sensor fusion by introducing deliberate temporal misalignment in one or more sensors.
+The attacker aims to disrupt the coherent integration of multi-modal sensor data, thereby inducing errors in perception.
+The resultant temporal inconsistencies can lead to incorrect perception results, particularly in safety-critical applications such as AD.
+Hence, the attacker’s key objective is to disrupt the perception to deteriorate the downstream tasks and, eventually, to compromise the safety and integrity of the control decision.
+III-C Proposed DejaVu Attack
+In this part, we elaborate on our proposed temporal misalignment attack, named DejaVu attack, where an adversary maliciously introduces a timing delay $\delta_{i}^{\text{mal}}$ to one or more sensors with a goal of creating misaligned sensor fusion and creating a false perception of the surroundings.
+Under DejaVu attack, the attacker compromises the network to add an additional malicious delay of $\delta_{i}^{\text{mal}}$.
+Hence, the attacker transmits a compromised data packet $\big(\tilde{x}_{i}^{(\tilde{t}_{i})},t_{i}\big)$ to the fusion ECU with the outdated semantic content $\tilde{x}_{i}^{(\tilde{t}_{i})}$ which was captured at global time $\tilde{t}_{i}$ instead of $t_{i}$, but with the updated time stamp $t_{i}$.
+In this case, $t_{i}=\tilde{t}_{i}+\delta_{i}^{\text{mal}}$, and the attacker delayed that transmission by $\delta_{i}^{\text{mal}}$ either through (such as $\scriptstyle C_{1}$$\scriptstyle C_{3}$).
+Therefore, $t_{\text{sys}}$ can be expressed as:
+$t_{\text{sys}}=\underbrace{\tilde{t}_{i}+\delta_{i}^{\text{mal}}}_{\displaystyle{t_{i}}}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}}$ (2)
+III-C1 Distribution of Delay $\delta_{i}^{\text{mal}}$
+The malicious delay $\delta_{i}^{\text{mal}}$ can vary over time–depending on the attacker’s intent, and be crafted to cause maximum disruption in sensor fusion.
+In a benign scenario, there will be no temporal misalignment, and sensor data will be received and processed in a timely manner.
+However, the attacker can launch different forms of DejaVu attack by controlling $\delta_{i}^{\text{mal}}$ in different ways.
+For instance, in Table II, we provide two possible DejaVu attacks scenarios against $S_{i}$, which are elaborated below:
+TABLE II: Two Types of DejaVu Attack Strategy Against $S_{i}$
+Attack Name Attack Type Delay Distribution $\delta_{i}^{\text{mal}}$
+Constant Delay Constant Constant, $k$
+Random Delay Random $\text{Uniform}(0,k)$
+Constant Delay
+This attack strategy introduces a fixed delay of $k$ frames in the sensor data stream.
+The attacker can achieve this by creating desynchronization among the clocks, tampering with sensor timestamps, or delaying data at the communication layer.
+Impact: The fusion model still receives temporally consistent data but with a lag in the target modality.
+This can lead to delayed (misaligned) perception (i.e., missing or shifted bounding boxes) in critical applications like object detection, where real-time perception is essential.
+Random Delay
+In this attack strategy, each frame experiences a different delay, randomly sampled from the range $[0,k]$.
+This strategy not only disrupts the real-time requirement but also disrupts the temporal sequence of sensor data, which is crucial for object tracking.
+Impact: The fusion system struggles to maintain the proper ordering of sensor inputs, leading to degraded perception accuracy.
+This will cause erratic behavior in time-sensitive sequential applications, particularly for object tracking and autonomous navigation.
+III-C2 Impact of the DejaVu Attack
+Based on the attacker’s capabilities, she can compromise either only one modality, which we define as an unimodal DejaVu attack (Uni-DejaVu), or she can compromise multiple modalities together, defined as a multimodal DejaVu attack (Mul-DejaVu).
+Under the Uni-DejaVu attack with compromised, where only sensor $S_{1}$ is compromised, the compromised perception result $\tilde{\mathcal{Y}}^{(t_{\text{sys}})}$ at time $t_{\text{sys}}$ can be expressed as:
+$\tilde{\mathcal{Y}}^{(t_{\text{sys}})}=\mathcal{H}_{\theta}\big(\mathcal{F}_{\theta}\big(\tilde{f}_{1}^{(\tilde{t}_{1})},f_{2}^{({t}_{2})},\dots,f_{m}^{({t}_{m})}\big)\big)$ (3)
+Similarly, under the Mul-DejaVu attack with all the sensors compromised, the compromised perception result $\tilde{\mathcal{P}}(t_{\text{sys}})$ at time $t_{\text{sys}}$ can be expressed as:
+$\tilde{\mathcal{Y}}^{(t_{\text{sys}})}=\mathcal{H}_{\theta}\big(\mathcal{F}_{\theta}\big(\tilde{f}_{1}^{(\tilde{t}_{1})},\tilde{f}_{2}^{(\tilde{t}_{2})},\dots,\tilde{f}_{m}^{(\tilde{t}_{m})}\big)\big)$ (4)
+Under the DejaVu attack, the system will perceive the surroundings in a different way than that of the benign case $\tilde{\mathcal{Y}}^{(t_{\text{sys}})}\neq{\mathcal{Y}}^{(t_{\text{sys}})}$, which can lead to erroneous vehicle control decisions.
+IV Experimental Settings
+This section describes the experimental setup to assess the DejaVu attack across different MMF models and datasets.
+IV-A Datasets
+We evaluate our approach on two widely used multimodal AD datasets for 3D object detection and multi-object tracking: KITTI Tracking and NuScenes.
+[…]
+NuScenes Dataset.
+The NuScenes Dataset [31] was collected in Boston (USA) and Singapore, focusing on dense urban traffic under diverse conditions.
+It provides six surround-view RGB cameras, a Velodyne HDL-32E LiDAR (20 Hz), and five radars, with all annotations sampled at 2 Hz.
+The dataset consists of 1000 driving scenes, each 20 seconds long, with 3D bounding boxes and tracking IDs for 23 classes, including vehicles, pedestrians, bicycles, and traffic barriers.
+Figure 3: Uni-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
+Figure 4: Mul-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
+Figure 5: Uni-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
+IV-B Models
+To systematically evaluate the effect of DejaVu on MMF with different downstream tasks, we consider the following models:
+3D Object Detection: We evaluate DejaVu on two representative MMF-based 3D object detection models.
+i) MVXNet [28], trained on the KITTI dataset, is an early fusion-based architecture that projects LiDAR point clouds into pseudo-image space and fuses them with camera image features at the voxel level.
+ii) BEVFusion [29] is a more advanced architecture, trained on the NuScenes dataset, that unifies multi-modal sensor inputs in the bird’s eye view (BEV) representation space.
+BEVFusion is widely used in industry-grade AD software stacks, including Autoware.
+Multi-Object Tracking (MOT): For evaluating tracking performance under DejaVu, we consider MMF-JDT [32], trained on the KITTI tracking dataset, is a joint detection and tracking model that incorporates early and mid-level fusion strategies to align image and point cloud features for improved object association over time.
+Figure 6: Mul-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
+IV-C Attack Settings
+To evaluate the impact of DejaVu attack on MMF models, we introduce controlled delays in one or both sensor modalities (camera and LiDAR), as mentioned in Table II, and analyze the corresponding degradation in model performance.
+We systematically assess how different degrees and types of temporal misalignment affect the 3D object detection and multi-object tracking.
+We introduce both constant and random delays in one or both modalities, as defined in Table II.
+Specifically, we use a delay parameter $k\in\{0,1,2,3,4,5\}$, where $k=0$ represents a perfectly synchronized sensor, and $k=5$ denotes a maximum delay of five frames for the affected modality.
+We consider both Uni-DejaVu and Mul-DejaVu attacks.
+In Uni-DejaVu attack, delay is introduced in either the camera or the LiDAR input while keeping the other modality synchronized.
+In Mul-DejaVu attacks, both camera and LiDAR streams are delayed independently, leading to varying degrees of temporal misalignment.
+We use the pretrained weights for the target model provided with the official implementations.
+We focus exclusively on attacking the test dataset by applying the defined temporal delays.
+IV-D Evaluation Metrics
+To assess the impact of DejaVu, we analyze the model performance using task-specific evaluation metrics.
+For 3D object detection, we evaluate the models using mean average precision (mAP), which quantifies the accuracy of detected objects, as well as nuScenes Detection Score (NDS), particularly for the NuScenes dataset.
+For MOT, we use standard tracking metrics such as higher order tracking accuracy (HOTA), detection accuracy (DetA), association accuracy (AssA), multiple object tracking accuracy (MOTA), and Identity Switches (IDSW), which measure the effectiveness of object association across frames.
+IV-E Software Implementation
+We implement and evaluate DejaVu using Python 3.8 and PyTorch, utilizing open-source frameworks including MMDetection3D [59] and OpenPCDet [60].
+Experiments were conducted on a server running Ubuntu 20.04.6 LTS with an Intel Xeon Gold 5520 (16 cores, 2.20GHz), 128GB RAM, and three NVIDIA RTX 6000 Ada GPUs.
+V Evaluation Results
+This section presents the DejaVu attack impact on different MMF models and datasets, and discusses the key findings from the evaluation.
+V-A Impact of DejaVu Attack on 3D Object Detection
+We investigate the effectiveness of the proposed DejaVu attack on multimodal 3D object detection using MVXNet and BEVFusion, evaluated on the KITTI and nuScenes datasets, respectively.
+We analyze the detection performance of 3D object detection across different object classes under varying levels of unimodal and multimodal sensor delay.
+Object detection models do not process sequential information; instead, their performance is affected only by frame-wise delays in each modality at any particular time, whether the delays are constant or random.
+Therefore, for simplicity and consistency, we only analyze the attack under the constant delay setting.
+Figure 7: Overall Mul-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for all the object classes regarding i) NDS, ii) mAP.
+(a) Constant Delay Scenario
+(b) Random Delay Scenario
+Figure 8: DejaVu attack impacts on multi-object (car) tracking performance of MMF-JDT on KITTI tracking dataset with respect to different metrics.
+V-A1 MVXNet on KITTI Dataset
+Fig.
+3 presents the 3D object detection performance of MVXNet under Uni-DejaVu attacks, where either the camera or LiDAR input is delayed independently.
+Under benign (zero-delay) conditions, MVXNet achieves strong mAP across most object classes: approximately 84.1 for cars, 87.6 for pedestrians, and 73.2 for cyclists.
+The left plot shows that delaying the camera input alone–under Uni-DejaVu camera attacks—has minimal effect on performance across all classes, with nearly constant mAPs.
+In contrast, the right plot highlights the model’s high sensitivity to LiDAR delays: a 1-frame delay causes the car mAP to collapse from 84.1 to 9.7 ($\downarrow$88.5%), pedestrian mAP from 87.6 to 34.2 ($\downarrow$60.9%), and cyclist mAP from 73.2 to 32.9 ($\downarrow$55.1%), with further degradation as delay increases.
+This indicates that LiDAR data is significantly more critical than camera input in MVXNet’s perception pipeline; hence, MVXNet’s high vulnerability against Uni-DejaVu attack against LiDAR.
+However, Fig.
+4 shows the mAP heatmaps across combinations of camera and LiDAR delays under Mul-DejaVu attacks.
+Although the 1-frame LiDAR delay drops mAP from 55.1% to 88.5% for different objects, camera delay has almost no effect, indicating the dominance of LiDAR in MVXNet.
+{keyfinding}
+Key Findings.
+MVXNet heavily depends on LiDAR input for 3D object detection.
+Camera delay, under both Uni- or Mul-DejaVu attack, has almost no effect, but even minor LiDAR misalignment leads to severe performance degradation.
+When both modalities are delayed, the impact is still only dominated by the LiDAR stream.
+V-A2 BEVFusion on nuScenes Dataset
+Fig.
+5 shows BEVFusion’s 3D detection performance under Uni-DejaVu attacks.
+With no delay, the model achieves high mAP across most classes—approximately 88 for car and pedestrian, and slightly lower for the rest.
+With a 1-frame camera delay, mAP drops slightly across all classes (for instance, car mAP drops by 7.4%), but remains stable with further delays, showing BEVFusion is slightly vulnerable against Uni-DejaVu camera attacks.
+In contrast, introducing a 1-frame LiDAR delay results in a substantial reduction in mAP for specific object classes, with performance dropping by 65.6% for cars (from 88.8 to 30.5), 35.1% for pedestrians (from 88.1 to 57.2), and 38.6% for motorcycles (from 75.9 to 46.6).
+Although performance remains stable with further LiDAR delays, this finding shows BEVFusion’s higher vulnerability against Uni-DejaVu LiDAR attacks.
+Similarly, Fig.
+6 illustrates BEVFusion’s performance under Mul-DejaVu attacks for all the objects.
+Although delaying the camera alone has a minimal impact on performance, combining this with a delay in the LiDAR stream causes a significant drop in mAP.
+For instance, a 1-frame delay in the camera or LiDAR stream reduces car mAP by 7.4% and 65.5%, respectively.
+However, when both modalities are delayed simultaneously by one frame, the mAP drops dramatically by 89.2% (from 88.8 to 9.6), highlighting a substantial compounding effect.
+This trend is consistently observed across all object classes.
+Fig.
+7 further shows the impact of Mul-DejaVu on the performance of BEVFusion on the nuScenes dataset regarding average mAP and NDS score.
+{keyfinding}
+Key Findings.
+BEVFusion is slightly affected by camera delays but is highly affected by LiDAR delays, further underscoring LiDAR’s dominating role in 3D object detection.
+However, the impact becomes significant if both sensors are delayed simultaneously, even just by one frame.
+Figure 9: Schematic diagram of hardware-in-the-loop Automotive Ethernet testbed.
+Figure 10: ROS System time vs (top) Timestamp of Camera and LiDAR messages, and (bottom) Message content of Camera and LiDAR messages with
+Figure 11: Demonstration of replay attacks where malicious LiDAR messages closely follow genuine ones, increasing synchronization likelihood.
+(a) False negative scenario.
+(b) False positive scenario.
+Figure 12: Impact of delayed yet valid LiDAR data on Autoware.
+a) The ego misses the oncoming truck, causing a head-on collision, and b) the Ego vehicle brakes for a non-existent object from delayed data.
+V-B Impact of DejaVu Attack on Multi Object Tracking
+We investigate the effectiveness of the proposed DejaVu attack on MOT algorithm using MMF-JDT on the KITTI dataset.
+We analyze the tracking performance for cars under varying levels of delays under Uni- and Mul-DejaVu attack scenarios.
+V-B1 MMF-JDT on KITTI Dataset
+This part studies the impact of DejaVu attacks on MMF-JDT evaluated on the KITTI tracking dataset.
+For this evaluation, we consider both types of delays: constant (Fig.
+8a) and random (Fig.
+8b).
+Across both attack scenarios, the tracking performance declines as camera delay increases.
+Along with other tracking metrics, MOTA and IDSW suffer noticeable degradation as camera delays increase under Uni-DejaVu attacks.
+For instance, IDSW—a metric that captures identity switches and ideally should be low, increases dramatically under increasing camera delays, underscoring the disruption in tracking consistency caused by DejaVu attacks.
+Furthermore, the values of different metrics across the heatmaps of Fig.
+8a suggest that while camera delay under Uni-DejaVu deteriorates the performance, delaying both the camera and LiDAR by exactly the same delay (i.e., constant delay scenario) under Mul-DejaVu attack diminishes the attack impact and mostly retains the performance.
+On the other hand, heatmaps in Fig.
+8b show that Mul-DejaVu attacks with random delays remain effective as different delays in both modalities break the sequence, making object tracking considerably more difficult.
+For instance, under the Mul-DejaVu constant attack scenario with a five-frame delay, MOTA decreases by only 1.4%, whereas the random attack scenario results in a 60.5% drop.
+{keyfinding}
+Key Findings.
+In contrast to 3D object detection tasks, MOT appears to rely more heavily on camera inputs.
+This may be attributed to the fact that MOT does not require precise 3D bounding boxes; instead, the rich texture information in camera images may offer more effective contrastive representations than sparse point clouds.
+As a result, DejaVu attacks can substantially impair MOT performance, particularly under Uni-DejaVu (camera delay) and Mul-DejaVu with random delay scenarios.
+V-C Hardware-in-the-loop Testbed
+In this part, we present our hardware-in-the-loop experiment for DejaVu, conducted on an AE testbed.
+As illustrated in Fig.
+9, the testbed consists of three Raspberry Pi, representing camera, LiDAR, and fusion node.
+To ensure reproducibility and precise control over experimental conditions, we utilize KITTI Tracking Dataset for the respective camera and LiDAR node.
+To enable TSN and AE functionality, each Raspberry Pi is equipped with a RealTime TSN HAT and a media converter, and the nodes are connected via an AE switch.
+ROS2 serves as the middleware for data distribution (i.e., publishing/subscribing messages).
+Published messages contain timestamps and sensor content (.png file for camera and .bin file for LiDAR), which are sent sequentially at 1-frame per second.
+The fusion node subscribes to the camera and LiDAR topics and aligns the sensor messages according to their timestamps using ROS2’s $ApproximateTimeSynchronizer$ filtering utility.
+This utility maintains queues of incoming camera and LiDAR messages and aligns them based on their timestamps, permitting a slight delay between matched messages.
+The fusion node then performs MMF-based downstream processing on the aligned sensor data.
+DejaVu Attack Impact on AE Testbed.
+In this experiment, we assume that the attacker is targeting the LiDAR sensor with the attacker capability $\scriptstyle C_{2}$ (see Section III-B1).
+As shown in Fig.
+10(top), the attacker first publishes six LiDAR messages with benign timestamp; hence they are very close to the camera messages’ timestamps.
+The attacker then deliberately introduces a constant offset of 5 seconds to the timestamps of the next eleven LiDAR messages, while continuing to transmit genuine, real-time data.
+As a result, the timestamps of these LiDAR messages are abruptly increased, as depicted in the figure.
+In the Fig.
+10(bottom), the message content is accurately synchronized for camera and LiDAR pairs, for all benign timestamps, which are shown in green rectangles.
+When the attacker sends manipulated timestamps, the synchronizer forces camera messages to align with LiDAR messages inaccurately, which are shown using red arrows in the figure.
+This demonstrates that though the contents of camera and LiDAR messages were sequential, the synchronizer fails to align them accurately during the attack, as it prioritizes the timestamp of the messages.
+The impact of this attack is similar to the attack we see in Fig.
+1(c).
+Figure 13: Impact of delayed but valid LiDAR data on Autoware: the induced SLAM drift propagates to the planning module, ultimately causing the vehicle to collide with the curb.
+V-D Simulation-Based End-to-End AD Setup
+We conduct our experiments using Autoware, an open-source full-stack autonomous driving framework.
+Autoware is widely adopted in both commercial and public-sector deployments, including Level 4 autonomy trials and government-funded programs (e.g., the U.S.
+Department of Transportation’s CARMASM initiative).
+Its extensive use in real-world systems makes it a realistic and representative platform for evaluating the safety and robustness of autonomous driving pipelines.
+To simulate real-world driving scenarios in a reproducible and controllable setting, we integrate Autoware with AWSIM Lab—a Unity-based, open-source simulator developed as part of the Autoware ecosystem—that provides high-fidelity urban environments and realistic sensor simulation.
+The simulated vehicle is equipped with a representative sensor suite including GNSS, IMU, three Velodyne VLP-16 LiDARs, and a traffic light camera.
+ROS2 (Humble) works as the middleware, enabling seamless and real-time communication between AWSIM’s simulated sensors and Autoware full autonomy stack, allowing us to evaluate the impact of DejaVu attack in a safe yet realistic environment.
+Our experiments are conducted on Tokyo’s Nishishinjuku district road map.
+DejaVu Attack Impact on AD Simulation.
+We demonstrate the DejaVu attack in a full-stack autonomous driving pipeline, targeting the LiDAR sensor under attacker capability $\scriptstyle C_{3}$ (see Section III-B1).
+Given LiDAR’s dominant role in 3D perception, the attacker impersonates three legitimate LiDAR nodes by subscribing to their respective ROS2 topics and monitoring inter-frame intervals (10 Hz) to predict the next transmission times.
+The attacker stores recent point cloud messages and, at the time of attack, publishes forged messages with previously captured data with updated timestamps—just before the expected legitimate message.
+This increases the likelihood that the forged message is selected by the time-based synchronizer if it aligns more closely with the timestamps of other modalities (e.g., IMU, camera).
+Fig.
+11 illustrates how these delayed messages are positioned and transmitted on the same topics, effectively impersonating genuine LiDAR messages in real time.
+When the forged messages are utilized in the downstream tasks, in the most severe case, the system completely misses the presence of an actual oncoming vehicle, resulting in a head-on collision at an intersection—despite nearby objects being within sensor range (Fig.
+12a).
+This constitutes a false negative perception failure with life-threatening implications.
+In another scenario, delayed LiDAR data causes the ego vehicle to perceive a non-existent obstacle—a vehicle that has already passed—leading to unnecessary emergency braking (Fig.
+12b).
+This false positive event can create rear-end collision risks.
+In both cases, the outdated sensor data was used to repeatedly overwrite fresh messages, degrading the temporal integrity of the perception pipeline.
+Beyond object-level failures, we observe broader impacts across the autonomy stack.
+The tracker may assign separate IDs to genuine and delayed instances of the same object, interpreting them as distinct entities.
+Similarly, temporal inconsistencies introduced in the LiDAR stream desynchronize SLAM modules, leading to localization drift and control failures such as veering off-lane or collisions with curbs and roadside objects (as shown in Fig.
+13).
+VI Discussion
+VI-A Attack Limitations
+While the DejaVu attack demonstrates the vulnerability of multimodal perception systems to temporal misalignment, there are several limitations that constrain its applicability in real-world settings.
+First, the attack has not been validated on a physical vehicle platform, where additional practical challenges such as sensor noise, actuator delays, and system integration issues may affect both the feasibility and effectiveness of the attack.
+Besides, the attack assumes a high level of knowledge about the target system, including the sensors and network architecture.
+In practice, the attack requires adversarial access to the in-vehicle network.
+Although not trivial, prior work shows that physical access through the OBD-II port, remote exploitation of infotainment systems, or supply-chain insertion can provide such capabilities [18].
+However, the adversary would still need to perform an exploration phase to gather this information, which introduces additional complexity and may limit the ability to execute the attack stealthily.
+Moreover, the evaluation does not account for real-world network-induced delays and variability, which could impact the targeted timing manipulation strategies.
+In actual deployment, the presence of stochastic latency and jitter may reduce the precision with which an attacker can control temporal misalignment, potentially diminishing the attack’s effectiveness.
+Finally, the evaluation was conducted exclusively on an offline dataset with relatively low frame rates.
+Real-time systems typically operate at higher frame rates, and the metrics and thresholds used in this offline evaluation may not directly translate to real-world performance.
+Consequently, the practical impact of the attack on deployed autonomous systems may differ from the results observed in the experimental study.
+VI-B Defense Strategies Against DejaVu
+Hardening defenses can reduce the attack surface by securing sensor timestamps before fusion.
+This can be achieved via authenticated, hardware-anchored timestamps with cryptographic signatures, monotonic sequence numbers, and hardware-level clocks, such as network interface controller/SoC real-time clock (RTC).
+Combining multiple time sources—PTP, GNSS, and local RTCs—further strengthens temporal integrity.
+While these measures add overhead, they significantly raise the barrier to adversarial manipulation.
+Detection defenses can identify temporal misalignments in real time.
+Techniques include intermodality temporal-consistency analysis of embeddings, kinematic cross-checks using IMU, odometry, and other controller area networks (CAN) signals, and statistical monitoring of monotonicity, jitter, and freshness counters, before fusion.
+These mechanisms can detect out-of-order or replayed frames, allowing rapid response to potential attacks.
+Mitigation techniques can limit the impact of detected misalignments on vehicle control.
+Delay-aware adaptive fusion compensates for inter-modal lags, weighs sensor inputs, and computes confidence scores.
+If confidence is low, conservative measures, such as slowing down, increasing spacing, or lowering autonomy, can be applied, ensuring safety even under temporal attacks.
+VII Conclusion
+This work presents DejaVu, a temporal misalignment attack that exploits synchronization vulnerabilities in multimodal perception systems for autonomous driving.
+Through extensive evaluations on state-of-the-art 3D object detection and multi-object tracking models, we uncover modality-specific vulnerabilities: 3D detection models are predominantly reliant on LiDAR and suffer severe degradation—with up to 88.5% drop in mAP—from even a single-frame LiDAR delay, while MOT models exhibit heightened sensitivity to camera stream disruptions, with MOTA dropping by 73% under just three-frame camera delays.
+These findings highlight the critical need for synchronization-aware design in perception architectures and emphasize the importance of robust temporal consistency checks in safety-critical autonomous systems.
+LLM Usage Disclosure.
+LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.